### v5

# OpenStereo: A Comprehensive Benchmark for Stereo Matching and Strong Baseline

Abstract
Stereo matching aims to estimate the disparity between matching pixels in a stereo image pair, which is of great importance to robotics, autonomous driving, and other computer vision tasks. Despite the development of numerous impressive methods in recent years, replicating their results and determining the most suitable architecture for practical application remains challenging. Addressing this gap, our paper introduces a comprehensive benchmark focusing on practical applicability rather than solely on performance enhancement. Specifically, we develop a flexible and efficient stereo matching codebase, called OpenStereo. OpenStereo includes training and inference codes of more than 10 network models, making it, to our knowledge, the most complete stereo matching toolbox available. Based on OpenStereo, we conducted experiments and have achieved or surpassed the performance metrics reported in the original paper. Additionally, we carry out an exhaustive analysis and deconstruction of recent developments in stereo matching through comprehensive ablative experiments. These investigations inspired the creation of StereoBase, a strong baseline model. Our StereoBase ranks 1st on SceneFlow, KITTI 2015, 2012 (Reflective) among published methods and achieves the best performance across all metrics. In addition, StereoBase has strong cross-dataset generalization. Code is available at https://github.com/XiandaGuo/OpenStereo.
11footnotetext: Joint second authors.44footnotetext: Corresponding authors
Figure 1: Comparison with other state-of-the-art stereo methods [24, 40, 1, 46, 5, 13, 39, 12, 51] on SceneFlow [28], KITTI 2012 [11], 2015 [29] leaderboards.
1 Introduction
Stereo matching is a fundamental topic in the field of computer vision, aiming to compute the disparity between a pair of rectified stereo images. It plays a crucial role in numerous applications such as robotics [60], autonomous driving [30, 39], and augmented reality [53], as it enables depth perception and 3D reconstruction of the observed scene.
Traditional stereo matching algorithms typically rely on matching corresponding image regions between the left and right views based on their similarity measures. Several techniques have been proposed in the literature for stereo matching, including methods based on gray-level information [2, 23, 56], region-based approaches [62, 33], and energy optimization methods [37, 16]. Gray-level-based methods [2, 23, 56] compute disparities by matching pixels or groups of pixels with similar intensities in the left and right images. Region-based approaches [62, 33] group pixels into larger regions and then compute correspondences between these regions. Energy optimization methods [37, 16] formulate the stereo matching problem as an energy minimization problem, where the goal is to find the disparity map that minimizes a certain energy function. While these methods have achieved some success in stereo matching, they suffer from certain limitations and challenges, such as occlusions, textureless regions, and computational complexity.
Recently, with the support of large synthetic datasets [28, 36, 38, 11, 29], CNNs-based stereo matching methods [19, 5, 58, 13, 52, 25, 7, 46] has achieved impressive results. These methods have shown remarkable performance in terms of accuracy and efficiency. Unlike traditional methods, CNN-based stereo matching methods are end-to-end frameworks that take a pair of stereo images as input and directly predict the disparity between them. In addition to the accuracy and efficiency benefits, CNN-based stereo matching methods have shown the potential for generalization to new environments and scenes beyond the training data. This is mainly due to the ability of CNNs to learn feature representations from raw data that are invariant to scale, rotation, and illumination changes. Furthermore, the availability of large-scale datasets has enabled the use of deep learning techniques for stereo matching, allowing for more complex and sophisticated models to be trained.
Nevertheless, we find that not all datasets are accompanied by official evaluation tools. For example, the DrivingStereo [55] dataset does not provide specific evaluation scripts, making comparative assessments challenging. The SceneFlow [28] dataset, with its finalpass and cleanpass data varieties, complicates fair model comparisons. Generalization experiments for stereo matching algorithms typically train on the SceneFlow dataset and evaluate on KITTI2012 [11], KITTI2015 [29], ETH3D [38], and Middlebury [36]. However, the inconsistency in evaluation data volume and metric definitions across studies leads to comparative discrepancies. Furthermore, even when source codes for models are available, reproducing the reported benchmarks is exceedingly challenging, e.g. STTR [24]. Hence, there’s a pressing need for a comprehensive benchmark study within the stereo matching community to enhance practicality and ensure consistent comparisons. To achieve this objective, we introduce a versatile stereo matching codebase: OpenStereo.
To promote scalability and adaptability, OpenStereo offers the following features: (1) Modual design, researchers can define a new model without the need to alter the model code itself by simply modifying a yaml configuration file. (2)Various frameworks, including Concatenation-based [5, 58], Correlation-based [52, 45, 46, 1], Interlaced-based [39], Group-wise-correlation-based [51], Combine-based methods [13], and Difference-based [21]. (3)Various datasets, including SceneFlow [28], KITTI2012 [11], KITTI2015 [29], Middlebury [36], ETH3D [38] and DrivingStereo [55] dataset. (4) State-of-the-art methods, including PSMNet [5], GwcNet [13], AANet [52], FADNet++ [46], CFNet [40], STTR [24], CoEx [1], CascadeStereo [12], MobileStereoNet [39] and IGEV [51].
Leveraging OpenStereo, we rigorously reassess various officially stated conclusions by meticulously re-implementing the ablation studies, including data augmentation, backbone architectures, cost construction, disparity regression and refinement processes. Based on the insights gleaned from these ablation experiments, we introduce StereoBase, a model that sets a new benchmark, surpassing recently proposed methods in terms of performance. Furthermore, we undertake a thorough investigation to establish that StereoBase is not only powerful but also serves as an empirically robust baseline model for stereo matching, demonstrating exceptional efficacy and resilience across diverse testing scenarios.
In summary, this paper makes significant contributions to the field of stereo matching research from three key aspects. (1) OpenStereo, a unified and extensible platform, enables researchers to conduct comprehensive stereo matching studies. (2) With the support of OpenStereo, we conduct a profound revisitation and thorough deconstruction of recent stereo matching methodologies. (3) We introduce StereoBase, which sets a new benchmark with EPE of 0.34 on SceneFlow [28] and ranks 1st on KITTI2015 [29] and 2012(Reflective)[11] leaderboards among published methods. Furthermore, StereoBase exhibits exceptional cross-dataset generalization capabilities. These contributions collectively advance the state-of-the-art in stereo matching, offering both insights and practical tools to the stereo matching research community.
2 Related Work
2.1 Stereo Matching
With the rapid development of CNNs, significant progress has been made in stereo matching. Based on the network pipeline of stereo matching, stereo matching methods can be roughly grouped into two categories [46], including the encoder-decoder network with 2D convolution (ED-Conv2D) and the cost volume matching with 3D convolution (CVM-Conv3D).
Figure 2: Timeline of Stereo Matching Models. This figure illustrates the timeline of various stereo matching models that have been proposed in the literature. The top part shows ED-conv2D-based models, while the bottom part shows CVM-conv3D-based models. Each model is labeled with its name and authors.
Stereo Matching with CVM-Conv3D The CVM-Conv3D methods are proposed to improve the performance of depth estimation [19, 5, 13, 58, 59, 12, 7, 44, 1, 41, 39, 22, 26, 50, 31]. These methods learn disparities from a 4D cost volume, which is mainly constructed by concatenating left feature maps with their corresponding right counterparts across each disparity level [5]. GCNet [19] firstly introduced a novel approach that combines 3D encoder-decoder architecture with a 2D convolutional network to obtain a dense feature representation, which is used to regularize a 4D concatenation volume. Following GCNet [19], PSMNet [5] proposes an approach for regularizing the concatenation volume by leveraging a stacked hourglass 3D convolutional neural network in tandem with intermediate supervision. In order to enhance the expressiveness of the cost volume and ultimately improve performance in ambiguous regions, GwcNet [58] proposes the group-wise correlation volume and ACVNet [50] proposes the attention concatenation volume. GANet [58] proposes a semi-global aggregation layer and local guided aggregation layer to capture local and the whole-image cost dependencies respectively. Based on GANet [58], DSMNet [59] construct a robust “domain normalization” method to overcome the challenges in cross-domain generalization. CoEx [1] proposes a novel approach called Guided Cost volume Excitation (GCE), which leverages image guidance to construct a simple channel excitation of the cost volume. IGEV-Stereo [51] leverages an iterative geometry encoding volume to capture both local and non-local geometry information, which outperforms existing methods on KITTI benchmarks and achieves cross-dataset generalization and high inference efficiency. Besides, as a potential solution to reduce the need for humans to design a good volumetric method, Neural Architecture Search (NAS) [15, 27, 3] has been employed in stereo matching [7].
However, these CVM-Conv3D methods still suffer from low time efficiency and high memory requirements, which are far from real-time inference even on server GPUs. Therefore, it is essential to address the accuracy and efficiency problems for real-world applications.
Stereo Matching with ED-Conv2D The ED-Conv2D methods [57, 28, 21, 34, 54, 9, 35, 46, 52, 43, 40, 24, 25, 39, 61], which adopt networks with 2D convolutions to predict disparity, has been driven by the need for improved accuracy, computational efficiency, and real-time performance. One of the early deep learning-based stereo matching methods, MC-CNN (Matching Cost CNN) [57], was proposed to learn a matching cost function for improving performance in the cost aggregation and optimization stages. Subsequently, Mayer et al. [28] present end-to-end networks for the estimation of disparity, called DispNet, which is pure 2D CNN architectures. However, the model still faces challenges in capturing the matching features, resulting in poor estimation results. To overcome this challenge, the correlation layer is introduced in the end-to-end architecture [28, 8, 17, 18] to better capture the relationship between the left and right images. By incorporating this layer, the accuracy of the model is significantly improved. Furthermore, FADNet++ [46] proposes an innovative approach to efficient disparity refinement using residual learning [14] in a coarse-to-fine manner. AutoDispNet [35] applied neural architecture search to automatically design stereo matching network structures. More recently, Croco-Stereo [47] shows that large-scale pre-training can be successful for stereo matching through well-adapted pretext tasks. This method can achieve state-of-the-art performance without using task-specific designs, like correlation volume or iterative estimation.
These works represent the significant progress that has been made in the field of stereo matching, highlighting the diverse range of methods and architectures that have been proposed to address the challenges associated with this problem.
2.2 Codebase
Numerous infrastructure code platforms have been developed in the deep learning research community, with the aim of facilitating research in specific fields. One such platform is OpenGait [10], a gait recognition library. OpenGait thoroughly examines the latest advancements in gait recognition, providing novel perspectives for subsequent research in this domain. In object detection, MMDetection [6] and Detectron2 [49] have emerged as an all-encompassing resource for several favored detection techniques. In pose estimation, OpenPose [4] has developed the first open-source system that operates in real-time for detecting the 2D pose of multiple individuals, including the detection of key points for the body, feet, hands, and face. In stereo matching, it is noteworthy that not all datasets are accompanied by official evaluation tools. For instance, the DrivingStereo [55] dataset does not have official evaluation codes, and there is a lack of unified tools for assessing the model generalization across different domains. This absence of standardized evaluation resources contributes to the observed discrepancies in cross-domain evaluations of the same model as reported in different studies. Therefore, it is time to build a benchmark for stereo matching.
3 OpenStereo
In recent years, there has been a proliferation of new frameworks and evaluation datasets for stereo matching. However, the lack of a unified and fair evaluation platform in this field is a significant issue that cannot be ignored. To address this challenge and promote academic research and practical application we have developed OpenStereo, a pyTorch-based [32] toolbox that provides a reliable and standardized evaluation framework for stereo matching.
Figure 3: The design principles of proposed codebase OpenStereo.
3.1 Design Principles of OpenStereo
As shown in Figure 3, our developed OpenStereo covers the following highlight features.
Modular Design. OpenStereo adopts a modular design, greatly facilitating researchers in exploring new networks. By simply modifying a YAML configuration file, researchers can define a new model without the need to alter the model code itself. This design significantly lowers the barriers for researchers to extend or integrate additional algorithms and modules within the framework. This approach empowers researchers to freely combine and customize their algorithms with minimal code composition, enhancing the framework’s usability and adaptability.
Compatibility with various frameworks. Currently, more and more stereo matching methods have emerged, such as Concatenation-based [19, 5, 58, 7], Correlation-based [52, 45, 46, 1], Group-wise-correlation-based [51], Difference-based [21], Interlaced-based [39], and Combine-based methods [13, 40]. As mentioned above, many open-source methods have a narrow focus on their specific models, making it challenging to extend to multiple frameworks. However, OpenStereo provides a solution to this problem by supporting all of the aforementioned frameworks. With OpenStereo, researchers and practitioners can easily compare and evaluate different stereo matching models under a standardized evaluation protocol.
Support for various evaluation datasets. OpenStereo is a comprehensive tool that includes datasets commonly used by researchers in stereo matching. OpenStereo not only supports synthetic stereo datasets such as SceneFlow [28], but also five real-world datasets: KITTI2012 [11], KITTI2015 [29], ETH3D [38], Middlebury [36], and DrivingStereo [55]. We introduce a suite of bespoke functions, meticulously crafted for each dataset, encompassing everything from initial data preprocessing to the final stages of evaluation. The evaluation module includes the submission of the results to all benchmarks.
Support for state-of-the-arts. In our work, we have successfully replicated various state-of-the-art stereo matching methods, including PSMNet [5], GwcNet [13], AANet [52], FADNet++ [46], CFNet [40], STTR [24], CoEx [1], CascadeStereo [12], MobileStereoNet [39] and IGEV [51]. Notably, the performance metrics we achieved, in most cases, surpass those reported in their original publications. This comprehensive replication effort provides a valuable resource for novices in the field, offering a rich array of official examples to facilitate an effective and efficient start in stereo matching research.
3.2 Main Modules
Technically, we follow the design of most PyTorch deep learning projects and divide OpenStereo into three modules, data, modeling, and evaluation.
Data module includes three key components: data loader, which is responsible for loading and preprocessing the data; data sampler, which is used to select a subset of data for each iteration; and data transform, which applies various transformations to the data to increase the diversity and complexity of the training set.
Modeling module is built upon the BaseModel base class, which predefines many behaviors of the stereo matching during both training and testing phases. This module consists of four essential components, namely the Feature Extraction, Cost Calculation, Cost Aggregation, and Disparity Processor, which are critical to current stereo matching algorithms.
Evaluation module is utilized to assess the performance of the obtained model. Recognizing that different datasets require distinct evaluation protocols, we have incorporated these varied protocols directly into OpenStereo, significantly simplifying the evaluation process.
4 Revisit Deep Stereo Matching
4.1 Recheck on Previous Methods
PSMNet [5] is one of the most crucial and frequently cited networks in stereo matching. The official implementation code is available at https://github.com/pytorch/vision. GwcNet [13] employs a Group-wise Correlation Stereo Network to exploit both local and global matching cost volumes (https://github.com/xy-guo/GwcNet). AANet [52] focuses on adaptively fusing multi-scale features through a lightweight and efficient adaptive aggregation module ( https://github.com/haofeixu/aanet). FADNet++ [46] employs a two-stage architecture, consisting of a feature-metric aggregation stage and a disparity refinement stage. (https://github.com/HKBU-HPML/FADNet). CFNet [40] utilizes a collaborative feature learning strategy to enhance the representation capacity of features in the matching cost computation (https://github.com/gallenszl/CFNet). STTR [24] focuses on geometric correspondence problems, including stereo matching, by employing transformers to capture long-range spatial and temporal context information (https://github.com/mli0603/stereo-transformer). CoEx [1] focuses on context-aware features and cost volume aggregation (https://github.com/antabangun/coex). CascadeStereo [12] adopts a cascaded architecture for accurate and efficient disparity estimation, whose official code is available at (https://github.com/alibaba/cascade-stereo). MobileStereoNet [39] builds upon the StereoNet [21], which proposes Interlacing Cost Volume Construction (https://github.com/cogsys-tuebingen/mobilestereonet). IGEV [51] builds a combined geometry encoding volume that not only encodes both geometric and contextual information but also intricately captures fine-grained details of local matching (https://github.com/gangweiX/IGEV).
Table 1: Quantitative evaluation on the SceneFlow [28] test. For each model, the specific category on the SceneFlow used is consistent with the original paper. Models marked with are OpenStereo implementations, others are from the original paper. Underline refers to evaluation in the non-occluded regions only [24]. Bold: Best
Method SceneFlow
TrainSize Categories EPE(px)
STTR [24] $256\times 512256 × 512$ FinalPass 0.43
STTR [24] $256\times 512256 × 512$ FinalPass 0.40
PSMNet [5] $256\times 512256 × 512$ CleanPass 1.09
PSMNet [5] $256\times 512256 × 512$ CleanPass 0.96
CFNet [40] $256\times 512256 × 512$ FinalPass 1.04
CFNet [40] $256\times 512256 × 512$ FinalPass 0.96
AANet [52] $288\times 576288 × 576$ CleanPass 0.87
AANet [52] $288\times 576288 × 576$ CleanPass 0.80
Mobilestereo-2D [39] $256\times 512256 × 512$ FinalPass 1.14
Mobilestereo-2D [39] $256\times 512256 × 512$ FinalPass 1.10
Mobilestereo-3D [39] $256\times 512256 × 512$ FinalPass 0.80
Mobilestereo-3D [39] $256\times 512256 × 512$ FinalPass 0.79
GwcNet [13] $256\times 512256 × 512$ FinalPass 0.76
GwcNet [13] $256\times 512256 × 512$ FinalPass 0.74
COEX [1] $288\times 576288 × 576$ FinalPass 0.68
COEX [1] $288\times 576288 × 576$ FinalPass 0.67
FADNet++ [46] $384\times 768384 × 768$ CleanPass 0.76
FADNet++ [46] $384\times 768384 × 768$ CleanPass 0.65
CascadeStereo [12] $256\times 512256 × 512$ FinalPass 0.72
CascadeStereo [12] $256\times 512256 × 512$ FinalPass 0.67
IGEV [51] $320\times 736320 × 736$ FinalPass 0.47
IGEV [51] $320\times 736320 × 736$ FinalPass 0.46
Table 2: Quantitative evaluation on KITTI2015 leadboard.
Method KITTI2015
D1-bg D1-fg D1-all
PSMNet [5] 1.86 4.62 2.32
PSMNet [5] 1.80 4.58 2.26
IGEV [51] 1.38 2.67 1.59
IGEV [51] 1.44 2.31 1.59
For benchmarking, it is critical to ensure that the results are reliable and trustworthy. To achieve this, we conducted our experiments using SceneFlow [28] and KITTI2015 [29] datasets. As shown in Table 1, in most cases, the reproduced performances of OpenStero are better than the results reported by the original papers. Regarding the KITTI2015 dataset, submission constraints led us to limit our leaderboard contributions to reproductions of the widely recognized PSMNet [5] and the latest state-of-the-art IGEV [51], as demonstrated in Table 2. We stand by the integrity of our benchmark, asserting its role in providing a reliable and valid evaluation of stereo matching models. OpenStereo is designed to offer the research community in stereo matching a standardized, comprehensive platform for method assessment. This facility enables meaningful and comparative analyses across various models, thereby serving as an valuable tool for the innovation and scrutiny of emerging algorithms.
4.2 Denoising Stereo Matching
With the support of OpenStereo, a comprehensive reevaluation of various stereo matching methods is conducted, including data augmentation, feature extraction, cost construction, disparity prediction and refinement.
Table 3: The model undergoes training and evaluation on the finalpass of the SceneFlow [28] training and test sets. KITTI2015 [29] training set, consisting of 200 images, is only employed to evaluate the generalizability of models under various data augmentation conditions. RC stands for RandomCrop. CES represents ColorAug, Erase, and Scale. HFlip denotes both images of a stereo and disparity are horizontally flipped. HSFlip horizontally flips both images in the stereo pair and then swaps them. VFlip involves vertically flipping both images in the stereo pair along with the disparity, inverting their top-bottom orientation. denotes the method proposed in this paper.
Data Augmentation SceneFlow KITTI15
EPE(px) EPE(px) D1_all
[…]
BatchRandom 0.5688 2.78 13.20
[…]
4.2.1 Data Augmentation
While five data augmentation techniques are prevalent in stereo matching [25, 51], including random crop, color augmentation, eraser transform, flip, and spatial transform, the empirical efficacy of each method specifically for stereo vision tasks remains an underexplored area in the field. Our work extends beyond these common practices, exploring not only these techniques but also experimenting with two additional methods: BatchRandom and Shift, to further understand their impact on stereo matching performance. BatchRandom denotes the technique applied during model training where the size of the random crop varies for each batch. The Shift involves randomly shifting the right image horizontally by a certain range and then cropping both images to a specified size. As shown in Table 3, most data augmentation strategies, except for the combined use of erase transform and random crop, lead to a decline in the model’s performance metrics of SceneFlow. This is because stereo matching involves pixel-level matching, and these data augmentations (color augmentation, flip, and spatial transform) can affect the alignment of pixels. The erase transform benefits stereo matching by enhancing the model’s capacity to cope with occlusions, missing data, and diverse real-world scenarios. It encourages robust feature learning and generalization, ultimately improving model performance. Furthermore, the combination of RandomCrop(320×736) with color augmentation delivers the most striking improvements in Kitti2015 dataset performance, achieving the lowest EPE of 1.61 px and the lowest D1_all of 7.81. Such improvements highlight the critical role of color augmentation in boosting the model’s resilience against the variability of color and lighting that naturally occurs in diverse driving environments.
Table 4: Ablation study of different backbones and pre-training methods on the finalpass of SceneFlow test datasets [28]. Flops and Params represent the computational complexity and parameters within the whole model, respectively.
Backbone Pre-train Flops Params EPE
FadNet++ [46] None 28.39G 1.77M 0.8633
MobilenetV2 100 [51] None 31.83G 2.78M 0.8227
PSMNet [5] None 130.87G 4.46M 0.8068
MobilenetV2 100 [51] None 31.83G 2.78M 0.8227
MobilenetV2 100 [51] Timm [48] 31.83G 2.78M 0.7132
MobilenetV2 120d [51] Timm [48] 38.75G 5.21M 0.6603
4.2.2 Feature Extraction
In Table 4, we delve into an ablation study to examine the impact of different backbones and pre-training methods. Initially, we evaluate different backbones without any pre-training. The backbone of FadNet++ [46], despite its low parameter count of 1.77M, results in an EPE of 0.8633, indicating less precision in disparity estimation. Employing MobilenetV2 100 [51] as a backbone improves the EPE to 0.8227, with a slight increase in computational demand (31.83G flops) and parameters (2.78M). The backbone of PSMNet [5], while the most computationally intensive (130.87G flops) with the highest number of parameters (4.46M), achieves a further reduction in EPE to 0.8068, suggesting that the complexity of the network can contribute to the enhancement of stereo matching performance. Furthermore, we delve into the impact of pre-training on the MobilenetV2 100 backbone. Without pre-training, the EPE remains at 0.8227. However, when pre-trained, we observe a substantial improvement, as the EPE drops to 0.7132. This underscores the value of leveraging large-scale image datasets to enhance feature extraction capabilities before fine-tuning specific tasks like stereo matching. In conclusion, our ablation study highlights the critical role of both the architecture’s complexity and pre-training in refining the stereo matching model’s accuracy.
Table 5: Ablation study results of cost volume on the finalpass of SceneFlow test datasets [28]. For these experiments, one-quarter features from stereo images are utilized to construct the cost volume. Gwc represents Group-wise correlation volume [13]. Cat stands for Concatenation volume [5]. G8-C16, G16-C24 and G32-C48 combine Group-wise correlation volume and Concatenation volume [13]. Channel represents the channel of cost volume. Dims refers to the dimensions of cost volume.
Cost Volume Dims Channel Flops Params EPE
Difference 3D - 17.44G 2.40M 1.02
Correlation 3D - 24.80G 4.01M 0.81
Interlaced8 4D 8 126.87G 2.83M 0.70
Gwc8 4D 8 31.83G 2.78M 0.71
Gwc16 4D 16 75.28G 3.89M 0.66
Gwc24 4D 24 147.53G 5.75M 0.63
Gwc32 4D 32 248.60G 8.34M 0.62
Gwc48 4D 48 537.15G 15.73M 0.60
Cat24 4D 24 148.36G 5.78M 0.65
Cat48 4D 48 537.99G 15.76M 0.61
Cat64 4D 64 941.79G 26.09M 0.60
G8-C16 4D 24 148.36G 5.78M 0.62
G16-C24 4D 40 379.30G 11.69M 0.60
G32-C48 4D 80 1460.80G 39.37M 0.60
4.2.3 Cost Construction
In Table 5, an ablation study on various cost volume strategies for stereo matching is presented. The study begins with simpler 3D cost volume methods like Difference [21] and Correlation [46], yielding higher EPE of 1.02 and 0.81, respectively, at lower computational costs. This suggests that while efficient, these methods may lack the nuanced disparity capture necessary for complex scenes. The Interlaced8 model, introduced by MobileStereoNet [39], achieves the same EPE comparable to the Gwc8 model. However, its computational expense is substantially higher, with a flop count of 126.87G which is significantly larger than that of the Gwc8 model. The group-wise correlation and concatenation models demonstrate a clear trend: as the channel depth increases, the EPE improves, indicating improved disparity estimations through richer feature capture. The combined volume (G8-C16) offers a more optimal balance between computational load and disparity estimation accuracy, which achieves an EPE of 0.62. G16-C24 and G32-C48 do not result in significant improvement in EPE, which stabilizes at 0.60, despite a dramatic increase in computational load, especially for G32-C48 which demands 1460.80 Gflops and has 39.37M parameters. These results highlight the delicate balance between accuracy and computational efficiency in designing cost volumes for disparity estimation. While deeper and combined volumes reduce the EPE, the gains might be marginal compared to the significant increase in computational requirements, raising questions about the practicality of these approaches in resource-constrained environments.
Table 6: Ablation study results of disparity regression and refinement on the finalpass of SceneFlow test datasets [28]. ArgMin refers to Differentiable ArMin. Context stands for ContextUpasmple [42, 25, 51].
Regression Refinement Flops Params EPE
ArgMin None 26.37G 2.69M 0.76
ArgMin RGBRefine [52] 53.15G 2.81M 0.72
ArgMin Context 31.83G 2.78M 0.71
ArgMin Context+RGBRefine [52] 58.61G 2.89M 0.69
ArgMin Context+DRNetRefine [52] 58.57G 2.89M 0.69
ArgMin ConvGRU [25, 51] 1363.71G 12.51M 0.46
4.2.4 Disparity Regression and Refinement
The Differentiable ArgMin [20, 5, 13, 40, 50, 12, 58, 39, 46] introduced by GCNet [20], calculates initial disparity by converting matching costs into probabilities via softmax and then computing a weighted sum of these probabilities across all disparity levels. As shown in Table 6, in this ablation study on disparity refinement for the SceneFlow test datasets, various strategies show differing impacts on model performance. Without refinement, the model has an EPE of 0.76. RGBRefine and Context methods slightly improve EPE to 0.72 and 0.71, respectively, with a modest increase in computational resources. Combining these methods further reduces EPE to 0.69, indicating marginal benefits from their integration. However, employing ConvGRU refinement leads to a substantial improvement, reducing EPE to 0.46, albeit at a significant cost in computational complexity (1363.71 Gflops) and model size (12.51M). This highlights a trade-off between accuracy improvements and increased computational demands.
4.3 Analysis and Discussion
4.3.1 Necessity of Comprehensive Ablation Study
In the evolving landscape of deep stereo matching, comprehensive ablation studies play a pivotal role in deciphering the effectiveness of different components and strategies. A thorough ablation study goes beyond mere performance metrics; it uncovers the underlying mechanics of different algorithms, revealing their strengths and weaknesses in various scenarios. For instance, different data augmentation techniques may yield contrasting effects on the model’s ability to match stereo images accurately. Similarly, the impact of varying backbones, cost volume configurations, and disparity regression methods on the overall performance can be profound. Understanding the specific contributions of each component is crucial for building more efficient and effective stereo matching systems.
4.3.2 Necessity of A Strong Baseline
A strong baseline in deep stereo matching research is critical for several key reasons. First, it serves as a vital reference point, enabling a clear assessment of new methods against an established standard. Second, a strong baseline allows for precise evaluation of the impact of specific changes, whether they are new data augmentation methods, different network architectures, or innovative disparity estimation techniques. This helps in isolating and understanding the contribution of each component to the overall performance. Additionally, a solid baseline ensures fair and meaningful comparisons across studies, providing a common ground for evaluating different research outcomes. This is crucial for maintaining consistency and validity in comparative analyses. In summary, a strong baseline is essential for meaningful advancements in deep stereo matching, ensuring that new developments are substantial, accurately assessed, and broadly applicable.
5 A Strong Pipeline: StereoBase
In light of our comprehensive analysis, the goal of this section is to establish a strong baseline model that surpasses existing standards in performance. StereoBase, our proposed solution, embodies this objective.
5.1 Pipeline
We build a strong baseline: StereoBase. StereoBase begins with a carefully curated data preprocessing step. For SceneFlow, KITTI2012, and KITTI2015, we implement an optimized mix of data augmentation techniques, including RandomCrop, ColorAug, and Erase Transform. This step ensures that the input data is not only varied and representative of real-world scenarios but also tailored to enhance stereo matching performance. Given the left and the right images, the pre-trained MobileNetV2 [48] networks are used as our foundational backbone, extracting features at a reduced scale of 1/4th the original size to form the cost volume. The cost volume utilizes a combined volume, which combines group-wise correlation volume with an 8-channel and concatenation volume with 16 channels, to ensure an optimal balance between computational load and disparity estimation accuracy. Hourglass networks [51] were implemented for cost aggregation, while convGRU [51] strategies were applied for the final disparity regression.
Table 7: Results on SceneFlow [28], KITTI 2012 [11], KITTI 2015 [29] leaderboard, and DrivingStereo [55]. All results on DrivingStereo [55] are derived using the OpenStereo. Underline refers to evaluation in the non-occluded regions only. Bold: Best.
Method SceneFlow KITTI 2012 KITTI 2015 DrivingStereo
EPE 3-noc 3-all D1-bg D1-fg D1-all EPE D1-all
STTR [24] 0.43 - - 1.70 3.61 2.01 OOM OOM
CFNet [40] 1.04 1.23 1.58 1.54 3.56 1.88 0.98 1.46
AANet [52] 0.87 1.91 2.42 1.65 3.96 2.03 2.91 15.16
COEX  [1] 0.68 1.55 1.93 1.74 3.41 2.02 1.34 2.70
FADNet++  [46] 0.76 - - 1.99 3.18 2.19 1.44 5.15
PSMNet [5] 1.09 1.49 1.89 1.86 4.62 2.32 1.19 2.26
GwcNet [13] 0.76 1.32 1.70 1.74 3.93 2.11 0.99 1.36
Mobilestereo-3D [39] 0.80 - - 1.75 3.87 2.10 1.06 1.61
CascadeStereo [12] 0.72 - - 1.59 4.03 2.00 1.31 2.84
IGEV-Stereo [51] 0.47 1.12 1.44 1.38 2.67 1.59 1.06 1.50
StereoBase(Ours) 0.34 1.00 1.26 1.28 2.26 1.44 1.15 2.19
5.2 Comparison with State-of-the-art
In our comprehensive evaluation, we benchmarked StereoBase against current state-of-the-art methods on SceneFlow, KITTI2012, 2015, and DrivingStereo. On the SceneFlow test set, we achieve a new SOTA EPE of 0.34. The quantitative comparisons, as summarized in Table 7, clearly illustrate the edge of StereoBase in handling complex stereo matching scenarios with greater precision. Further, we submitted our results to the KITTI2012 and 2015 leaderboard, where StereoBase outperformed all published methods across all metrics. On KITTI2015, our StereoBase outperforms IGEV [51] by 9.43% on D1-all metric, respectively. In addition, we evaluate the generalization performance of StereoBase. As shown in Table 8, StereoBase exhibited exceptional performance in a zero-shot setting, achieving state-of-the-art results. This evaluation further validates the adaptability and potential of StereoBase in handling diverse and challenging stereo vision tasks.
Table 8: Cross-domain evaluation on Middlebury, ETH3D, and KITTI all training sets. All methods are only trained on the Scene Flow datatest. Middlebury is tested on half-resolution. The model with indicates the implementation of OpenStereo. Bold: Best.
Method KITTI2012 KITTI2015 Middlebury ETH3D
D1-all(%) EPE D1-all(%) EPE bad 2.0(%) bad 1.0(%)
STTR 49.72 6.80 40.26 6.16 OOM 38.89
PSMNet 30.51 4.68 32.15 5.99 33.53 18.02
CFNet 13.64 2.27 12.09 2.89 23.91 7.67
AANet 7.23 1.27 7.72 1.41 22.45 18.77
Mobilestereo-2D 18.34 2.45 21.21 2.78 34.04 13.89
Mobilestereo-3D 18.96 2.79 19.69 3.40 29.32 13.71
GwcNet 23.05 2.76 25.19 3.58 29.87 14.54
COEX 12.08 1.80 11.00 2.48 25.17 11.43
FADNet++ 11.31 1.77 13.23 2.97 24.17 25.53
CascadeStereo 11.83 1.83 12.03 2.69 27.27 11.68
IGEV 7.80 1.40 6.35 1.31 10.51 4.44
StereoBase 4.85 0.99 5.35 1.18 9.76 3.12
6 Conclusion
This paper introduces OpenStereo, a codebase designed for deep stereo matching. Our initial endeavor involved the re-implementation of the most state-of-the-art methods within the OpenStereo framework. This comprehensive tool facilitates the extensive reevaluation of various aspects of stereo matching methodologies. Drawing on the insights gained from our exhaustive ablation studies, we proposed StereoBase. StereoBase not only demonstrates the capabilities of our platform but also sets a new standard in the field for future research and development. Through OpenStereo and StereoBase, we aim to contribute a substantial and versatile resource to the stereo matching community, fostering innovation and facilitating more effective and efficient research.

### v6

@@
 Abstract
-Stereo matching aims to estimate the disparity between matching pixels in a stereo image pair, which is of great importance to robotics, autonomous driving, and other computer vision tasks.
-Despite the development of numerous impressive methods in recent years, replicating their results and determining the most suitable architecture for practical application remains challenging.
-Addressing this gap, our paper introduces a comprehensive benchmark focusing on practical applicability rather than solely on performance enhancement.
+Stereo matching aims to estimate the disparity between matching pixels in a stereo image pair, which is important to robotics, autonomous driving, and other computer vision tasks.
+Despite the development of numerous impressive methods in recent years, determining the most suitable architecture for practical application remains challenging.
+Addressing this gap, our paper introduces a comprehensive benchmark focusing on practical applicability rather than solely on individual models for optimized performance.
 Specifically, we develop a flexible and efficient stereo matching codebase, called OpenStereo.
@@
 Based on OpenStereo, we conducted experiments and have achieved or surpassed the performance metrics reported in the original paper.
-Additionally, we carry out an exhaustive analysis and deconstruction of recent developments in stereo matching through comprehensive ablative experiments.
+Additionally, we conduct an exhaustive analysis and deconstruction of recent developments in stereo matching through comprehensive ablative experiments.
 These investigations inspired the creation of StereoBase, a strong baseline model.
-Our StereoBase ranks 1st on SceneFlow, KITTI 2015, 2012 (Reflective) among published methods and achieves the best performance across all metrics.
+Our StereoBase ranks 1st on SceneFlow, KITTI 2015, 2012 (Reflective) among published methods and performs best across all metrics.
 In addition, StereoBase has strong cross-dataset generalization.
 Code is available at https://github.com/XiandaGuo/OpenStereo.
-11footnotetext: Joint second authors.44footnotetext: Corresponding authors
-Figure 1: Comparison with other state-of-the-art stereo methods [24, 40, 1, 46, 5, 13, 39, 12, 51] on SceneFlow [28], KITTI 2012 [11], 2015 [29] leaderboards.
+11footnotetext: Joint second authors.22footnotetext: Corresponding authors:long.chen@ia.ac.cn
 1 Introduction
@@
 It plays a crucial role in numerous applications such as robotics [72], autonomous driving [37, 46], and augmented reality [64], as it enables depth perception and 3D reconstruction of the observed scene.
-Traditional stereo matching algorithms typically rely on matching corresponding image regions between the left and right views based on their similarity measures.
+Traditional stereo-matching algorithms typically rely on matching corresponding image regions between the left and right views based on their similarity measures.
 Several techniques have been proposed in the literature for stereo matching, including methods based on gray-level information [3, 29, 68], region-based approaches [73, 39], and energy optimization methods [44, 21].
-Gray-level-based methods [2, 23, 56] compute disparities by matching pixels or groups of pixels with similar intensities in the left and right images.
-Region-based approaches [62, 33] group pixels into larger regions and then compute correspondences between these regions.
-Energy optimization methods [37, 16] formulate the stereo matching problem as an energy minimization problem, where the goal is to find the disparity map that minimizes a certain energy function.
-While these methods have achieved some success in stereo matching, they suffer from certain limitations and challenges, such as occlusions, textureless regions, and computational complexity.
-Recently, with the support of large synthetic datasets [28, 36, 38, 11, 29], CNNs-based stereo matching methods [19, 5, 58, 13, 52, 25, 7, 46] has achieved impressive results.
-These methods have shown remarkable performance in terms of accuracy and efficiency.
-Unlike traditional methods, CNN-based stereo matching methods are end-to-end frameworks that take a pair of stereo images as input and directly predict the disparity between them.
-In addition to the accuracy and efficiency benefits, CNN-based stereo matching methods have shown the potential for generalization to new environments and scenes beyond the training data.
-This is mainly due to the ability of CNNs to learn feature representations from raw data that are invariant to scale, rotation, and illumination changes.
-Furthermore, the availability of large-scale datasets has enabled the use of deep learning techniques for stereo matching, allowing for more complex and sophisticated models to be trained.
-Nevertheless, we find that not all datasets are accompanied by official evaluation tools.
+With the support of large synthetic datasets [35, 43, 45, 14, 36, 66], CNNs-based stereo matching methods [24, 6, 70, 18, 61, 33, 10, 52, 60, 53, 8, 30, 32] has achieved impressive results.
+As shown in Figure 1, based on the network pipeline of stereo matching, CNNs-based stereo matching methods can be roughly grouped into two categories [52], including the encoder-decoder network with 2D convolution (ED-Conv2D) [35, 40, 65, 41, 51, 52, 61, 49, 31, 33, 46, 28, 56, 30, 16] and the cost volume matching with 3D convolution (CVM-Conv3D) [24, 6, 67, 55, 70, 18, 71, 12, 74, 15, 1, 10, 2, 47, 59, 60, 54, 9, 62].
+However, we find that different studies often employ varying data augmentation strategies, learning rates, learning rate optimization methods, and backbone architectures.
+This inconsistency makes it difficult to evaluate and compare various methods’ performance accurately.
+This inconsistency in experimental setups and methodologies makes it difficult to derive conclusive insights and hampers the objective assessment of advancements in stereo matching.
+Without a standardized benchmark, researchers struggle to identify the true impact of individual components and innovations.
+There is a lack of clear conclusions and exploration regarding data augmentation strategies, backbone selection, and cost construction methods in stereo matching.
+Moreover, not all datasets are accompanied by official evaluation tools.
 For example, the DrivingStereo [66] dataset does not provide specific evaluation scripts, making comparative assessments challenging.
@@
 Generalization experiments for stereo matching algorithms typically train on the SceneFlow dataset and evaluate on KITTI2012 [14], KITTI2015 [36], ETH3D [45], and Middlebury [43].
-However, the inconsistency in evaluation data volume and metric definitions across studies leads to comparative discrepancies.
-Furthermore, even when source codes for models are available, reproducing the reported benchmarks is exceedingly challenging, e.g.
-STTR [24].
-Hence, there’s a pressing need for a comprehensive benchmark study within the stereo matching community to enhance practicality and ensure consistent comparisons.
-To achieve this objective, we introduce a versatile stereo matching codebase: OpenStereo.
+However, due to the absence of a standard protocol for generalization experiments, different papers may yield inconsistent results for the same method.
+For instance, discussions on the generalization performance of IGEV [60] across works [60, 53, 17] exemplify this issue.
+Figure 1: Timeline of Stereo Matching Models. The top part shows ED-conv2D-based models, while the bottom part shows CVM-conv3D-based models. Each model is labeled with its name and authors.
+Hence, there’s a pressing need for a comprehensive benchmark study within the stereo-matching community to enhance practicality and ensure consistent comparisons.
+To achieve this objective, we introduce a versatile stereo-matching codebase: OpenStereo.
 To promote scalability and adaptability, OpenStereo offers the following features: (1) Modual design, researchers can define a new model without the need to alter the model code itself by simply modifying a yaml configuration file.
@@
 (4) State-of-the-art methods, including PSMNet [6], GwcNet [18], AANet [61], FADNet++ [52], CFNet [47], STTR [31], CoEx [2], CascadeStereo [15], MobileStereoNet [46] and IGEV [60].
-Leveraging OpenStereo, we rigorously reassess various officially stated conclusions by meticulously re-implementing the ablation studies, including data augmentation, backbone architectures, cost construction, disparity regression and refinement processes.
+Leveraging OpenStereo, we rigorously reassess various officially stated conclusions by meticulously re-implementing the ablation studies, including data augmentation, backbone architectures, cost construction, disparity regression, and refinement processes.
 Based on the insights gleaned from these ablation experiments, we introduce StereoBase, a model that sets a new benchmark, surpassing recently proposed methods in terms of performance.
-Furthermore, we undertake a thorough investigation to establish that StereoBase is not only powerful but also serves as an empirically robust baseline model for stereo matching, demonstrating exceptional efficacy and resilience across diverse testing scenarios.
-In summary, this paper makes significant contributions to the field of stereo matching research from three key aspects.
-(1) OpenStereo, a unified and extensible platform, enables researchers to conduct comprehensive stereo matching studies.
-(2) With the support of OpenStereo, we conduct a profound revisitation and thorough deconstruction of recent stereo matching methodologies.
-(3) We introduce StereoBase, which sets a new benchmark with EPE of 0.34 on SceneFlow [28] and ranks 1st on KITTI2015 [29] and 2012(Reflective)[11] leaderboards among published methods.
-Furthermore, StereoBase exhibits exceptional cross-dataset generalization capabilities.
-These contributions collectively advance the state-of-the-art in stereo matching, offering both insights and practical tools to the stereo matching research community.
+StereoBase is powerful and serves as an empirically robust baseline model for stereo matching, demonstrating exceptional efficacy and resilience across diverse testing scenarios.
+Our contributions are summarized as follows:
+  • $\bullet$
+    We introduce OpenStereo, a unified and extensible platform, which enables researchers to conduct comprehensive stereo matching studies.
+  • $\bullet$
+    We conduct a profound revisitation and thorough deconstruction of recent stereo-matching methodologies.
+  • $\bullet$
+    We introduce StereoBase, which sets a new benchmark with EPE of 0.34 on SceneFlow [35] and ranks 1st on KITTI2015 [36] and 2012(Reflective)[14] leaderboards among published methods.
 2 Related Work
@@
 Based on the network pipeline of stereo matching, stereo matching methods can be roughly grouped into two categories [52], including the encoder-decoder network with 2D convolution (ED-Conv2D) and the cost volume matching with 3D convolution (CVM-Conv3D).
-Figure 2: Timeline of Stereo Matching Models.
-This figure illustrates the timeline of various stereo matching models that have been proposed in the literature.
-The top part shows ED-conv2D-based models, while the bottom part shows CVM-conv3D-based models.
-Each model is labeled with its name and authors.
-Stereo Matching with CVM-Conv3D The CVM-Conv3D methods are proposed to improve the performance of depth estimation [19, 5, 13, 58, 59, 12, 7, 44, 1, 41, 39, 22, 26, 50, 31].
-These methods learn disparities from a 4D cost volume, which is mainly constructed by concatenating left feature maps with their corresponding right counterparts across each disparity level [5].
+Stereo Matching with CVM-Conv3D The CVM-Conv3D methods are proposed to improve the performance of depth estimation [24, 6, 67, 55, 70, 18, 71, 12, 74, 15, 1, 10, 2, 47, 59, 60, 54, 9, 62].
+These methods learn disparities from a 4D cost volume, mainly constructed by concatenating left feature maps with their corresponding right counterparts across each disparity level [6].
 GCNet [24] firstly introduced a novel approach that combines 3D encoder-decoder architecture with a 2D convolutional network to obtain a dense feature representation, which is used to regularize a 4D concatenation volume.
 Following GCNet [24], PSMNet [6] proposes an approach for regularizing the concatenation volume by leveraging a stacked hourglass 3D convolutional neural network in tandem with intermediate supervision.
-In order to enhance the expressiveness of the cost volume and ultimately improve performance in ambiguous regions, GwcNet [58] proposes the group-wise correlation volume and ACVNet [50] proposes the attention concatenation volume.
+To enhance the expressiveness of the cost volume and ultimately improve performance in ambiguous regions, GwcNet [70] proposes the group-wise correlation volume and ACVNet [59] proposes the attention concatenation volume.
 GANet [70] proposes a semi-global aggregation layer and local guided aggregation layer to capture local and the whole-image cost dependencies respectively.
-Based on GANet [58], DSMNet [59] construct a robust “domain normalization” method to overcome the challenges in cross-domain generalization.
+Based on GANet [70], DSMNet [71] constructs a robust “domain normalization” method to overcome the challenges in cross-domain generalization.
 CoEx [2] proposes a novel approach called Guided Cost volume Excitation (GCE), which leverages image guidance to construct a simple channel excitation of the cost volume.
-IGEV-Stereo [51] leverages an iterative geometry encoding volume to capture both local and non-local geometry information, which outperforms existing methods on KITTI benchmarks and achieves cross-dataset generalization and high inference efficiency.
+IGEV-Stereo [60] leverages an iterative geometry encoding volume to capture local and non-local geometry information, outperforming existing methods on KITTI benchmarks and achieving cross-dataset generalization and high inference efficiency.
 Besides, as a potential solution to reduce the need for humans to design a good volumetric method, Neural Architecture Search (NAS) [20, 34, 4] has been employed in stereo matching [10].
@@
 One of the early deep learning-based stereo matching methods, MC-CNN (Matching Cost CNN) [69], was proposed to learn a matching cost function for improving performance in the cost aggregation and optimization stages.
-Subsequently, Mayer et al.
-[28] present end-to-end networks for the estimation of disparity, called DispNet, which is pure 2D CNN architectures.
+Subsequently, Mayer et al [35] present end-to-end networks for the estimation of disparity, called DispNet, which is pure 2D CNN architectures.
 However, the model still faces challenges in capturing the matching features, resulting in poor estimation results.
@@
 Support for various evaluation datasets.
-OpenStereo is a comprehensive tool that includes datasets commonly used by researchers in stereo matching.
-OpenStereo not only supports synthetic stereo datasets such as SceneFlow [28], but also five real-world datasets: KITTI2012 [11], KITTI2015 [29], ETH3D [38], Middlebury [36], and DrivingStereo [55].
+OpenStereo is a comprehensive tool that not only supports synthetic stereo datasets such as SceneFlow [35], but also five real-world datasets: KITTI2012 [14], KITTI2015 [36], ETH3D [45], Middlebury [43], and DrivingStereo [66].
 We introduce a suite of bespoke functions, meticulously crafted for each dataset, encompassing everything from initial data preprocessing to the final stages of evaluation.
-The evaluation module includes the submission of the results to all benchmarks.
+The evaluation module includes the submission of the results to KITTI2012 [14] and KITTI2015 [36] leadboards.
 Support for state-of-the-arts.
 In our work, we have successfully replicated various state-of-the-art stereo matching methods, including PSMNet [6], GwcNet [18], AANet [61], FADNet++ [52], CFNet [47], STTR [31], CoEx [2], CascadeStereo [15], MobileStereoNet [46] and IGEV [60].
-Notably, the performance metrics we achieved, in most cases, surpass those reported in their original publications.
-This comprehensive replication effort provides a valuable resource for novices in the field, offering a rich array of official examples to facilitate an effective and efficient start in stereo matching research.
+As shown in Figure 3, the performance metrics we achieved, in most cases, surpass those reported in their original publications.
 3.2 Main Modules
@@
 4.1 Recheck on Previous Methods
-PSMNet [5] is one of the most crucial and frequently cited networks in stereo matching.
-The official implementation code is available at https://github.com/pytorch/vision.
-GwcNet [13] employs a Group-wise Correlation Stereo Network to exploit both local and global matching cost volumes (https://github.com/xy-guo/GwcNet).
-AANet [52] focuses on adaptively fusing multi-scale features through a lightweight and efficient adaptive aggregation module ( https://github.com/haofeixu/aanet).
-FADNet++ [46] employs a two-stage architecture, consisting of a feature-metric aggregation stage and a disparity refinement stage.
-(https://github.com/HKBU-HPML/FADNet).
-CFNet [40] utilizes a collaborative feature learning strategy to enhance the representation capacity of features in the matching cost computation (https://github.com/gallenszl/CFNet).
-STTR [24] focuses on geometric correspondence problems, including stereo matching, by employing transformers to capture long-range spatial and temporal context information (https://github.com/mli0603/stereo-transformer).
-CoEx [1] focuses on context-aware features and cost volume aggregation (https://github.com/antabangun/coex).
-CascadeStereo [12] adopts a cascaded architecture for accurate and efficient disparity estimation, whose official code is available at (https://github.com/alibaba/cascade-stereo).
-MobileStereoNet [39] builds upon the StereoNet [21], which proposes Interlacing Cost Volume Construction (https://github.com/cogsys-tuebingen/mobilestereonet).
-IGEV [51] builds a combined geometry encoding volume that not only encodes both geometric and contextual information but also intricately captures fine-grained details of local matching (https://github.com/gangweiX/IGEV).
-Table 1: Quantitative evaluation on the SceneFlow [28] test.
+Figure 3: Quantitative evaluation on the SceneFlow [35] test and KITTI2015 [36] leadboard.
 For each model, the specific category on the SceneFlow used is consistent with the original paper.
-Models marked with are OpenStereo implementations, others are from the original paper.
-Underline refers to evaluation in the non-occluded regions only [24].
-Bold: Best
-Method SceneFlow
-TrainSize Categories EPE(px)
-STTR [24] $256\times 512256 × 512$ FinalPass 0.43
-STTR [24] $256\times 512256 × 512$ FinalPass 0.40
-PSMNet [5] $256\times 512256 × 512$ CleanPass 1.09
-PSMNet [5] $256\times 512256 × 512$ CleanPass 0.96
-CFNet [40] $256\times 512256 × 512$ FinalPass 1.04
-CFNet [40] $256\times 512256 × 512$ FinalPass 0.96
-AANet [52] $288\times 576288 × 576$ CleanPass 0.87
-AANet [52] $288\times 576288 × 576$ CleanPass 0.80
-Mobilestereo-2D [39] $256\times 512256 × 512$ FinalPass 1.14
-Mobilestereo-2D [39] $256\times 512256 × 512$ FinalPass 1.10
-Mobilestereo-3D [39] $256\times 512256 × 512$ FinalPass 0.80
-Mobilestereo-3D [39] $256\times 512256 × 512$ FinalPass 0.79
-GwcNet [13] $256\times 512256 × 512$ FinalPass 0.76
-GwcNet [13] $256\times 512256 × 512$ FinalPass 0.74
-COEX [1] $288\times 576288 × 576$ FinalPass 0.68
-COEX [1] $288\times 576288 × 576$ FinalPass 0.67
-FADNet++ [46] $384\times 768384 × 768$ CleanPass 0.76
-FADNet++ [46] $384\times 768384 × 768$ CleanPass 0.65
-CascadeStereo [12] $256\times 512256 × 512$ FinalPass 0.72
-CascadeStereo [12] $256\times 512256 × 512$ FinalPass 0.67
-IGEV [51] $320\times 736320 × 736$ FinalPass 0.47
-IGEV [51] $320\times 736320 × 736$ FinalPass 0.46
-Table 2: Quantitative evaluation on KITTI2015 leadboard.
-Method KITTI2015
-D1-bg D1-fg D1-all
-PSMNet [5] 1.86 4.62 2.32
-PSMNet [5] 1.80 4.58 2.26
-IGEV [51] 1.38 2.67 1.59
-IGEV [51] 1.44 2.31 1.59
+Underline refers to evaluation in the non-occluded regions only STTR [31].
 For benchmarking, it is critical to ensure that the results are reliable and trustworthy.
-To achieve this, we conducted our experiments using SceneFlow [28] and KITTI2015 [29] datasets.
-As shown in Table 1, in most cases, the reproduced performances of OpenStero are better than the results reported by the original papers.
-Regarding the KITTI2015 dataset, submission constraints led us to limit our leaderboard contributions to reproductions of the widely recognized PSMNet [5] and the latest state-of-the-art IGEV [51], as demonstrated in Table 2.
-We stand by the integrity of our benchmark, asserting its role in providing a reliable and valid evaluation of stereo matching models.
+To achieve this, we conducted our experiments on SceneFlow [35] and KITTI2015 [36] datasets.
+As shown in Figure 3, the reproduced performances of OpenStereo are better than the results reported by the original papers.
+Regarding the KITTI2015 dataset, submission constraints led us to limit our leaderboard contributions to reproductions of the widely recognized PSMNet [6] and the latest state-of-the-art IGEV [60].
 OpenStereo is designed to offer the research community in stereo matching a standardized, comprehensive platform for method assessment.
@@
 4.2 Denoising Stereo Matching
-With the support of OpenStereo, a comprehensive reevaluation of various stereo matching methods is conducted, including data augmentation, feature extraction, cost construction, disparity prediction and refinement.
+With the support of OpenStereo, a comprehensive reevaluation of various stereo-matching methods is conducted, including data augmentation, feature extraction, cost construction, disparity prediction, and refinement.
+Our ablation studies have revealed some new insights.
+Figure 4: The visualization of stereo images and disparity with different data augmentation.
+The blue box represents the area where random cropping occurs during training.
+Notably, when the views are horizontally flipped, the disparity are multiplied by 11-1- 1, making the disparity map appear pure white.
 Table 1: The model undergoes training and evaluation on the finalpass of the SceneFlow [35] training and test sets.
-KITTI2015 [29] training set, consisting of 200 images, is only employed to evaluate the generalizability of models under various data augmentation conditions.
+KITTI2015 [36] training set, consisting of 200 images, is only employed to evaluate the generalizability of models.
 RC stands for RandomCrop.
-CES represents ColorAug, Erase, and Scale.
 HFlip denotes both images of a stereo and disparity are horizontally flipped.
@@
 VFlip involves vertically flipping both images in the stereo pair along with the disparity, inverting their top-bottom orientation.
-denotes the method proposed in this paper.
-Data Augmentation SceneFlow KITTI15
+CES represents ColorAug, Erase, and Scale.
+Settings used in our final model are underlined.
+Data Augmentation LR_scheduler SceneFlow KITTI15
 EPE(px) EPE(px) D1_all
-[…]
-[…]
-BatchRandom 0.5688 2.78 13.20
 […]
@@
 […]
-4.2.1 Data Augmentation
-While five data augmentation techniques are prevalent in stereo matching [25, 51], including random crop, color augmentation, eraser transform, flip, and spatial transform, the empirical efficacy of each method specifically for stereo vision tasks remains an underexplored area in the field.
-Our work extends beyond these common practices, exploring not only these techniques but also experimenting with two additional methods: BatchRandom and Shift, to further understand their impact on stereo matching performance.
-BatchRandom denotes the technique applied during model training where the size of the random crop varies for each batch.
-The Shift involves randomly shifting the right image horizontally by a certain range and then cropping both images to a specified size.
-As shown in Table 3, most data augmentation strategies, except for the combined use of erase transform and random crop, lead to a decline in the model’s performance metrics of SceneFlow.
-This is because stereo matching involves pixel-level matching, and these data augmentations (color augmentation, flip, and spatial transform) can affect the alignment of pixels.
-The erase transform benefits stereo matching by enhancing the model’s capacity to cope with occlusions, missing data, and diverse real-world scenarios.
-It encourages robust feature learning and generalization, ultimately improving model performance.
-Furthermore, the combination of RandomCrop(320×736) with color augmentation delivers the most striking improvements in Kitti2015 dataset performance, achieving the lowest EPE of 1.61 px and the lowest D1_all of 7.81.
-Such improvements highlight the critical role of color augmentation in boosting the model’s resilience against the variability of color and lighting that naturally occurs in diverse driving environments.
-Table 4: Ablation study of different backbones and pre-training methods on the finalpass of SceneFlow test datasets [28].
+[…]
+[…]
+[…]
+[…]
+Table 2: Ablation study of different backbones on the finalpass of SceneFlow test datasets [35].
 Flops and Params represent the computational complexity and parameters within the whole model, respectively.
-Backbone Pre-train Flops Params EPE
-FadNet++ [46] None 28.39G 1.77M 0.8633
-MobilenetV2 100 [51] None 31.83G 2.78M 0.8227
-PSMNet [5] None 130.87G 4.46M 0.8068
-MobilenetV2 100 [51] None 31.83G 2.78M 0.8227
-MobilenetV2 100 [51] Timm [48] 31.83G 2.78M 0.7132
-MobilenetV2 120d [51] Timm [48] 38.75G 5.21M 0.6603
+FLOPs are calculated at a resolution of $544\times 960544 × 960$.
+Backbone Type Pretrain Flops Params EPE
+MobilenetV2 100 [42] CNN 70.58G 2.78M 0.7737
+MobilenetV2 100 [42] CNN 70.58G 2.78M 0.6155
+MobilenetV2 100 [42] CNN 70.58G 2.78M 0.6155
+MobilenetV2 120d [42] CNN 85.93G 5.21M 0.5573
+EfficientNetV2 [48] CNN 157.52G 24.92M 0.5207
+RepViT [50] Trans. 101.35G 5.64M 0.5858
+MPViT [27] CNN&Trans. 283.35G 13.33M 0.5113
+4.2.1 LR_scheduler and Data Augmentation
+As shown in Table 1, MultiStepLR yields an EPE of 0.6839, while OneCycleLR achieves a lower EPE of 0.6155.
+This substantial difference underscores the crucial role of selecting an appropriate learning rate scheduler for stereo matching.
+The superior performance of OneCycleLR indicates its potential to improve model accuracy and robustness, making it a preferable choice over MultiStepLR for training stereo-matching models.
+Although five data augmentation techniques—random crop, color augmentation, eraser transform, flip, and spatial transform—are commonly used in stereo matching [33, 60], their empirical efficacy specifically for stereo matching has not been thoroughly explored.
+This study investigated these data augmentation strategies to address this gap.
+Visualizations of these augmentations are provided in Figure 4 for further insight.
+As shown in Table 1, most data augmentation strategies, except for the combined use of HSFlip and random crop, lead to a decline in the model’s EPE metric of SceneFlow.
+This is because stereo matching involves pixel-level matching, and these data augmentations (color augmentation and spatial transform) can affect the alignment of pixels.
+The combination of ColorAug, Erase, and Scale (CES) shows the best generalization performance on KITTI2015, with the lowest EPE of 1.56 and D1_all of 7.64, although it increases the EPE on SceneFlow to 0.7240.
+These findings underscore the importance of selecting appropriate data augmentation methods to enhance model accuracy and robustness.
 4.2.2 Feature Extraction
-In Table 4, we delve into an ablation study to examine the impact of different backbones and pre-training methods.
-Initially, we evaluate different backbones without any pre-training.
-The backbone of FadNet++ [46], despite its low parameter count of 1.77M, results in an EPE of 0.8633, indicating less precision in disparity estimation.
-Employing MobilenetV2 100 [51] as a backbone improves the EPE to 0.8227, with a slight increase in computational demand (31.83G flops) and parameters (2.78M).
-The backbone of PSMNet [5], while the most computationally intensive (130.87G flops) with the highest number of parameters (4.46M), achieves a further reduction in EPE to 0.8068, suggesting that the complexity of the network can contribute to the enhancement of stereo matching performance.
-Furthermore, we delve into the impact of pre-training on the MobilenetV2 100 backbone.
-Without pre-training, the EPE remains at 0.8227.
-However, when pre-trained, we observe a substantial improvement, as the EPE drops to 0.7132.
-This underscores the value of leveraging large-scale image datasets to enhance feature extraction capabilities before fine-tuning specific tasks like stereo matching.
-In conclusion, our ablation study highlights the critical role of both the architecture’s complexity and pre-training in refining the stereo matching model’s accuracy.
-Table 5: Ablation study results of cost volume on the finalpass of SceneFlow test datasets [28].
-For these experiments, one-quarter features from stereo images are utilized to construct the cost volume.
+As shown in Table 2, pretraining the backbone is crucial for stereo matching as it enhances the model’s ability to extract robust and informative features.
+Furthermore, the choice of backbone significantly influences the model’s performance and computational efficiency.
+MobilenetV2 [42] and EfficientNetV2 [48] are lightweight CNNs that are particularly efficient in extracting local features, which are crucial for stereo matching.
+Their designs allow them to perform well with relatively low computational complexity.
+RepViT [50] is a Transformer-based architecture, which excels in capturing long-range dependencies and global context.
+While RepViT captures global features well, it might struggle with the fine-grained, pixel-level accuracy required for precise stereo matching.
+MPViT [27] combines the strengths of both CNNs and Transformers.
+The CNN components effectively capture local features, while the Transformer components excel in modeling global context.
+This hybrid approach allows MPViT to leverage the advantages of both architectures, resulting in the lowest EPE.
+In summary, MobilenetV2 [42] offers a good balance for applications with limited computational resources, while more complex architectures like EfficientNetV2 and MPViT provide superior accuracy at the cost of higher computational requirements.
+To the best of our knowledge, our work is the first to explore the transformer-based feature extraction and the combination of CNN and transformer feature extraction for stereo matching.
+Table 3: Ablation study of cost volume on the finalpass of SceneFlow test datasets [35].
 Gwc represents Group-wise correlation volume [18].
 Cat stands for Concatenation volume [6].
-G8-C16, G16-C24 and G32-C48 combine Group-wise correlation volume and Concatenation volume [13].
-Channel represents the channel of cost volume.
-Dims refers to the dimensions of cost volume.
+G8-C16, G16-C24, and G32-C48 combine Gwc volume and Cat volume [18].
+Channel and Dims represent the channel and dimensions of the cost volume, respectively.
+FLOPs are calculated at a resolution of $544\times 960544 × 960$.
 Cost Volume Dims Channel Flops Params EPE
-Difference 3D - 17.44G 2.40M 1.02
-Correlation 3D - 24.80G 4.01M 0.81
-Interlaced8 4D 8 126.87G 2.83M 0.70
-Gwc8 4D 8 31.83G 2.78M 0.71
-Gwc16 4D 16 75.28G 3.89M 0.66
-Gwc24 4D 24 147.53G 5.75M 0.63
-Gwc32 4D 32 248.60G 8.34M 0.62
-Gwc48 4D 48 537.15G 15.73M 0.60
-Cat24 4D 24 148.36G 5.78M 0.65
-Cat48 4D 48 537.99G 15.76M 0.61
-Cat64 4D 64 941.79G 26.09M 0.60
-G8-C16 4D 24 148.36G 5.78M 0.62
-G16-C24 4D 40 379.30G 11.69M 0.60
-G32-C48 4D 80 1460.80G 39.37M 0.60
+Difference 3D - 38.68G 2.40M 1.02
+Correlation 3D - 54.99G 4.01M 0.81
+Interlaced8 4D 8 288.52G 2.83M 0.70
+Gwc8 4D 8 70.58G 2.78M 0.72
+Gwc16 4D 16 166.92G 3.89M 0.66
+Gwc24 4D 24 327.13G 5.75M 0.63
+Gwc32 4D 32 551.23G 8.34M 0.62
+Gwc48 4D 48 1191.07G 15.73M 0.60
+Cat24 4D 24 328.97G 5.78M 0.65
+Cat48 4D 48 1192.93G 15.76M 0.61
+Cat64 4D 64 2088.31G 26.09M 0.60
+G8-C16 4D 24 328.96G 5.78M 0.62
+G16-C24 4D 40 841.05G 11.69M 0.60
+G32-C48 4D 80 3239.17G 39.37M 0.60
 4.2.3 Cost Construction
 In Table 3, an ablation study on various cost volume strategies for stereo matching is presented.
-The study begins with simpler 3D cost volume methods like Difference [21] and Correlation [46], yielding higher EPE of 1.02 and 0.81, respectively, at lower computational costs.
+For these experiments, one-quarter of stereo image features are used to construct the cost volume.
+The study begins with simpler 3D cost volume methods: Difference [26] and Correlation [52], yielding higher EPE of 1.02 and 0.81, respectively, at lower computational costs.
 This suggests that while efficient, these methods may lack the nuanced disparity capture necessary for complex scenes.
 The Interlaced8 model, introduced by MobileStereoNet [46], achieves the same EPE comparable to the Gwc8 model.
-However, its computational expense is substantially higher, with a flop count of 126.87G which is significantly larger than that of the Gwc8 model.
+However, its computational expense is substantially higher, with a flop count of 288.52G, significantly larger than that of the Gwc8 model.
 The group-wise correlation and concatenation models demonstrate a clear trend: as the channel depth increases, the EPE improves, indicating improved disparity estimations through richer feature capture.
 The combined volume (G8-C16) offers a more optimal balance between computational load and disparity estimation accuracy, which achieves an EPE of 0.62.
-G16-C24 and G32-C48 do not result in significant improvement in EPE, which stabilizes at 0.60, despite a dramatic increase in computational load, especially for G32-C48 which demands 1460.80 Gflops and has 39.37M parameters.
+G16-C24 and G32-C48 do not significantly improve EPE, despite a dramatic increase in computational load, especially for G32-C48, which demands 3239.17Gflops and has 39.37M parameters.
 These results highlight the delicate balance between accuracy and computational efficiency in designing cost volumes for disparity estimation.
 While deeper and combined volumes reduce the EPE, the gains might be marginal compared to the significant increase in computational requirements, raising questions about the practicality of these approaches in resource-constrained environments.
-Table 6: Ablation study results of disparity regression and refinement on the finalpass of SceneFlow test datasets [28].
-ArgMin refers to Differentiable ArMin.
-Context stands for ContextUpasmple [42, 25, 51].
+Table 4: Ablation study of disparity regression and refinement on the finalpass of SceneFlow test datasets [35]. ArgMin refers to Differentiable ArMin. Context stands for ContextUpasmple [33, 60].
 Regression Refinement Flops Params EPE
-ArgMin None 26.37G 2.69M 0.76
-ArgMin RGBRefine [52] 53.15G 2.81M 0.72
-ArgMin Context 31.83G 2.78M 0.71
-ArgMin Context+RGBRefine [52] 58.61G 2.89M 0.69
-ArgMin Context+DRNetRefine [52] 58.57G 2.89M 0.69
-ArgMin ConvGRU [25, 51] 1363.71G 12.51M 0.46
+ArgMin None 58.47G 2.69M 0.76
+ArgMin RGBRefine [61] 117.85G 2.81M 0.72
+ArgMin Context 70.58G 2.78M 0.71
+ArgMin Context+RGBRefine [61] 129.95G 2.89M 0.69
+ArgMin Context+DRNetRefine [61] 129.88G 2.89M 0.69
+ArgMin ConvGRU [33, 60] 3023.88G 12.51M 0.46
 4.2.4 Disparity Regression and Refinement
 The Differentiable ArgMin [25, 6, 18, 47, 59, 15, 70, 46, 52] introduced by GCNet [25], calculates initial disparity by converting matching costs into probabilities via softmax and then computing a weighted sum of these probabilities across all disparity levels.
-As shown in Table 6, in this ablation study on disparity refinement for the SceneFlow test datasets, various strategies show differing impacts on model performance.
+As shown in Table 4, various strategies show differing impacts on model performance in this ablation study on disparity refinement for the SceneFlow test datasets.
 Without refinement, the model has an EPE of 0.76.
@@
 Combining these methods further reduces EPE to 0.69, indicating marginal benefits from their integration.
-However, employing ConvGRU refinement leads to a substantial improvement, reducing EPE to 0.46, albeit at a significant cost in computational complexity (1363.71 Gflops) and model size (12.51M).
+However, employing ConvGRU refinement substantially improves EPE 0.46, albeit at a significant cost in computational complexity (3023.88 Gflops) and model size (12.51M).
 This highlights a trade-off between accuracy improvements and increased computational demands.
@@
 Similarly, the impact of varying backbones, cost volume configurations, and disparity regression methods on the overall performance can be profound.
-Understanding the specific contributions of each component is crucial for building more efficient and effective stereo matching systems.
+Understanding the specific contributions of each component is crucial for building more efficient and effective stereo-matching systems.
 4.3.2 Necessity of A Strong Baseline
-A strong baseline in deep stereo matching research is critical for several key reasons.
+A strong baseline in deep stereo-matching research is critical for several key reasons.
 First, it serves as a vital reference point, enabling a clear assessment of new methods against an established standard.
@@
 5 A Strong Pipeline: StereoBase
+5.1 Pipeline
 In light of our comprehensive analysis, the goal of this section is to establish a strong baseline model that surpasses existing standards in performance.
-StereoBase, our proposed solution, embodies this objective.
-5.1 Pipeline
-We build a strong baseline: StereoBase.
-StereoBase begins with a carefully curated data preprocessing step.
-For SceneFlow, KITTI2012, and KITTI2015, we implement an optimized mix of data augmentation techniques, including RandomCrop, ColorAug, and Erase Transform.
-This step ensures that the input data is not only varied and representative of real-world scenarios but also tailored to enhance stereo matching performance.
-Given the left and the right images, the pre-trained MobileNetV2 [48] networks are used as our foundational backbone, extracting features at a reduced scale of 1/4th the original size to form the cost volume.
-The cost volume utilizes a combined volume, which combines group-wise correlation volume with an 8-channel and concatenation volume with 16 channels, to ensure an optimal balance between computational load and disparity estimation accuracy.
+StereoBase embodies this objective.
+As shown in Figure 5, given the left and the right images, the pre-trained MobileNetV2 [57] networks are used as our foundational backbone, extracting features at a reduced scale of 1/4th the original size to form the cost volume.
+The G8-C16 cost volume is utilized to achieve an optimal balance between computational load and disparity estimation accuracy.
 Hourglass networks [60] were implemented for cost aggregation, while convGRU [60] strategies were applied for the final disparity regression.
+Figure 5: Overview of our proposed StereoBase. GwcVolume represents Group-wise correlation volume [18]. CGEV refers to Combined Geometry Encoding Volume [60].
 Table 5: Results on SceneFlow [35], KITTI 2012 [14], KITTI 2015 [36] leaderboard, and DrivingStereo [66].
 All results on DrivingStereo [66] are derived using the OpenStereo.
-Underline refers to evaluation in the non-occluded regions only.
+Underline refers to evaluation in the non-occluded regions only STTR [31].
 Bold: Best.
@@
 STTR [31] 0.43 - - 1.70 3.61 2.01 OOM OOM
+PSMNet [6] 1.09 1.49 1.89 1.86 4.62 2.32 1.19 2.26
+GwcNet [18] 0.76 1.32 1.70 1.74 3.93 2.11 0.99 1.36
 CFNet [47] 1.04 1.23 1.58 1.54 3.56 1.88 0.98 1.46
 AANet [61] 0.87 1.91 2.42 1.65 3.96 2.03 2.91 15.16
+Mobilestereo-3D [46] 0.80 - - 1.75 3.87 2.10 1.06 1.61
 COEX  [2] 0.68 1.55 1.93 1.74 3.41 2.02 1.34 2.70
 FADNet++  [52] 0.76 - - 1.99 3.18 2.19 1.44 5.15
-PSMNet [5] 1.09 1.49 1.89 1.86 4.62 2.32 1.19 2.26
-GwcNet [13] 0.76 1.32 1.70 1.74 3.93 2.11 0.99 1.36
-Mobilestereo-3D [39] 0.80 - - 1.75 3.87 2.10 1.06 1.61
 CascadeStereo [15] 0.72 - - 1.59 4.03 2.00 1.31 2.84
 IGEV-Stereo [60] 0.47 1.12 1.44 1.38 2.67 1.59 1.06 1.50
+GANet+ADL [63] 0.50 0.98 1.29 1.38 2.38 1.55 - -
+NMRF-Stereo [17] 0.45 1.01 1.35 1.28 3.13 1.59 - -
+Selective-IGEV [53] 0.44 1.07 1.38 1.33 2.61 1.55 - -
+MoCha-Stereo [8] 0.41 1.06 1.35 1.36 2.43 1.53 - -
 StereoBase(Ours) 0.34 1.00 1.26 1.28 2.26 1.44 1.15 2.19
-5.2 Comparison with State-of-the-art
-In our comprehensive evaluation, we benchmarked StereoBase against current state-of-the-art methods on SceneFlow, KITTI2012, 2015, and DrivingStereo.
-On the SceneFlow test set, we achieve a new SOTA EPE of 0.34.
-The quantitative comparisons, as summarized in Table 7, clearly illustrate the edge of StereoBase in handling complex stereo matching scenarios with greater precision.
-Further, we submitted our results to the KITTI2012 and 2015 leaderboard, where StereoBase outperformed all published methods across all metrics.
-On KITTI2015, our StereoBase outperforms IGEV [51] by 9.43% on D1-all metric, respectively.
+5.2 Comparison with State-of-the-art Methods
+In our comprehensive evaluation, we benchmarked StereoBase against current state-of-the-art methods on SceneFlow [35], KITTI2012 [14], 2015 [36], and DrivingStereo [66].
+On the SceneFlow [35] test set, we achieve a new SOTA EPE of 0.34.
+The quantitative comparisons, as summarized in Table 5, clearly illustrate the edge of StereoBase in handling complex stereo-matching scenarios with greater precision.
+Further, we submitted our results to the KITTI2012 [14] and 2015 [36] leaderboards, where StereoBase outperformed all published methods across all metrics.
+On KITTI2015 [36], our StereoBase outperforms IGEV [60] by 9.43% on D1-all metric, respectively.
 In addition, we evaluate the generalization performance of StereoBase.
-As shown in Table 8, StereoBase exhibited exceptional performance in a zero-shot setting, achieving state-of-the-art results.
+As shown in Table 6, StereoBase exhibited exceptional performance in a zero-shot setting.
 This evaluation further validates the adaptability and potential of StereoBase in handling diverse and challenging stereo vision tasks.
@@
 CascadeStereo 11.83 1.83 12.03 2.69 27.27 11.68
-IGEV 7.80 1.40 6.35 1.31 10.51 4.44
+IGEV 4.88 0.98 5.16 1.18 8.47 3.53
 StereoBase 4.85 0.99 5.35 1.18 9.76 3.12
+Figure 6: Visualization results on KITTI2015 dataset.
 6 Conclusion
-This paper introduces OpenStereo, a codebase designed for deep stereo matching.
-Our initial endeavor involved the re-implementation of the most state-of-the-art methods within the OpenStereo framework.
-This comprehensive tool facilitates the extensive reevaluation of various aspects of stereo matching methodologies.
+This paper introduces OpenStereo, a benchmark designed for deep stereo matching.
+Our initial endeavor involved re-implementing the most state-of-the-art methods within the OpenStereo framework.
+This comprehensive tool facilitates the extensive reevaluation of various aspects of stereo-matching methodologies.
 Drawing on the insights gained from our exhaustive ablation studies, we proposed StereoBase.
+Our StereoBase ranks 1st on SceneFlow, KITTI 2015, 2012 (Reflective) among published methods and performs best across all metrics.
+In addition, StereoBase has strong cross-dataset generalization.
 StereoBase not only demonstrates the capabilities of our platform but also sets a new standard in the field for future research and development.
-Through OpenStereo and StereoBase, we aim to contribute a substantial and versatile resource to the stereo matching community, fostering innovation and facilitating more effective and efficient research.
+Through OpenStereo and StereoBase, we aim to contribute a substantial and versatile resource to the stereo-matching community, fostering innovation and facilitating more effective and efficient research.
+Acknowledgements.
+This work was supported by the National Natural Science Foundation of China under Grant 62373356 and the Open Projects Program of State Key Laboratory of Multimodal Artificial Intelligence Systems.

### v7

@@
 Despite the development of numerous impressive methods in recent years, determining the most suitable architecture for practical application remains challenging.
-Addressing this gap, our paper introduces a comprehensive benchmark focusing on practical applicability rather than solely on individual models for optimized performance.
+To address this gap, our paper introduces a comprehensive benchmark focusing on practical applicability rather than solely on individual models for optimized performance.
 Specifically, we develop a flexible and efficient stereo matching codebase, called OpenStereo.
@@
 It plays a crucial role in numerous applications such as robotics [66], autonomous driving [32, 41], and augmented reality [58], as it enables depth perception and 3D reconstruction of the observed scene.
-Traditional stereo-matching algorithms typically rely on matching corresponding image regions between the left and right views based on their similarity measures.
+Traditional stereo-matching algorithms typically match corresponding image regions between the left and right views based on their similarity measures.
 Several techniques have been proposed in the literature for stereo matching, including methods based on gray-level information [3, 26, 62], region-based approaches [67, 34], and energy optimization methods [38, 18].
-With the support of large synthetic datasets [35, 43, 45, 14, 36, 66], CNNs-based stereo matching methods [24, 6, 70, 18, 61, 33, 10, 52, 60, 53, 8, 30, 32] has achieved impressive results.
-As shown in Figure 1, based on the network pipeline of stereo matching, CNNs-based stereo matching methods can be roughly grouped into two categories [52], including the encoder-decoder network with 2D convolution (ED-Conv2D) [35, 40, 65, 41, 51, 52, 61, 49, 31, 33, 46, 28, 56, 30, 16] and the cost volume matching with 3D convolution (CVM-Conv3D) [24, 6, 67, 55, 70, 18, 71, 12, 74, 15, 1, 10, 2, 47, 59, 60, 54, 9, 62].
-However, we find that different studies often employ varying data augmentation strategies, learning rates, learning rate optimization methods, and backbone architectures.
+With the support of large synthetic datasets [30, 39, 40, 12, 31, 61], CNN-based stereo matching methods [21, 5, 15, 55] has achieved impressive results.
+As shown in Figure 1, based on the network pipeline of stereo matching, CNN-based stereo matching methods can be roughly grouped into two categories [48], including the encoder-decoder network with 2D convolution (ED-Conv2D) [30, 35, 60, 36, 48, 56, 45, 28, 29, 25, 51, 27] and the cost volume matching with 3D convolution (CVM-Conv3D) [21, 5, 64, 15, 65, 10, 68, 13, 1, 8, 2, 42, 54, 55, 7, 57].
+However, we find that different studies often employ various data augmentation strategies, learning rates, learning rate optimization methods, and backbone architectures.
 This inconsistency makes it difficult to evaluate and compare various methods’ performance accurately.
@@
 Moreover, not all datasets are accompanied by official evaluation tools.
-For example, the DrivingStereo [66] dataset does not provide specific evaluation scripts, making comparative assessments challenging.
 The SceneFlow [30] dataset, with its finalpass and cleanpass data varieties, complicates fair model comparisons.
@@
 To achieve this objective, we introduce a versatile stereo-matching codebase: OpenStereo.
-To promote scalability and adaptability, OpenStereo offers the following features: (1) Modual design, researchers can define a new model without the need to alter the model code itself by simply modifying a yaml configuration file.
+To promote scalability and adaptability, OpenStereo offers the following features: (1) Modulal design, researchers can define a new model without the need to alter the model code itself by simply modifying a YAML configuration file.
 (2)Various frameworks, including Concatenation-based [5, 64], Correlation-based [56, 47, 48, 2], Interlaced-based [41], Group-wise-correlation-based [55], Combine-based methods [15], and Difference-based [22].
-(3)Various datasets, including SceneFlow [35], KITTI2012 [14], KITTI2015 [36], Middlebury [43], ETH3D [45] and DrivingStereo [66] dataset.
+(3)Various datasets, including SceneFlow [30], KITTI2012 [12], KITTI2015 [31], Middlebury [39], and ETH3D [40].
 (4) State-of-the-art methods, including PSMNet [5], GwcNet [15], AANet [56], FADNet++ [48], CFNet [42], STTR [28], CoEx [2], CascadeStereo [13], MobileStereoNet [41] and IGEV [55].
-Leveraging OpenStereo, we rigorously reassess various officially stated conclusions by meticulously re-implementing the ablation studies, including data augmentation, backbone architectures, cost construction, disparity regression, and refinement processes.
+Leveraging OpenStereo, we rigorously reassess various officially stated conclusions by re-implementing the ablation studies, including data augmentation, backbone architectures, cost construction, disparity regression, and refinement processes.
 Based on the insights gleaned from these ablation experiments, we introduce StereoBase, a model that sets a new benchmark, surpassing recently proposed methods in terms of performance.
-StereoBase is powerful and serves as an empirically robust baseline model for stereo matching, demonstrating exceptional efficacy and resilience across diverse testing scenarios.
+StereoBase is powerful and serves as an empirically state-of-the-art (SOTA) baseline model for stereo matching, demonstrating exceptional efficacy and resilience across diverse testing scenarios.
 Our contributions are summarized as follows:
@@
     We introduce OpenStereo, a unified and extensible platform, which enables researchers to conduct comprehensive stereo matching studies.
+  • $\bullet$
+    We evaluate the impact of CNN, Transformer, and hybrid CNN-Transformer backbones on stereo matching. To our knowledge, this is the first work to explore these architectures.
   • $\bullet$
@@
 Based on the network pipeline of stereo matching, stereo matching methods can be roughly grouped into two categories [48], including the encoder-decoder network with 2D convolution (ED-Conv2D) and the cost volume matching with 3D convolution (CVM-Conv3D).
-Stereo Matching with CVM-Conv3D The CVM-Conv3D methods are proposed to improve the performance of depth estimation [24, 6, 67, 55, 70, 18, 71, 12, 74, 15, 1, 10, 2, 47, 59, 60, 54, 9, 62].
+Figure 2: The design principles of proposed codebase OpenStereo.
+Stereo Matching with CVM-Conv3D The CVM-Conv3D methods are proposed to improve the performance of depth estimation [21, 5, 59, 50, 64, 15, 65, 10, 68, 13, 1, 8, 2, 42, 54, 55, 57].
 These methods learn disparities from a 4D cost volume, mainly constructed by concatenating left feature maps with their corresponding right counterparts across each disparity level [5].
@@
 To enhance the expressiveness of the cost volume and ultimately improve performance in ambiguous regions, GwcNet [64] proposes the group-wise correlation volume and ACVNet [54] proposes the attention concatenation volume.
-GANet [70] proposes a semi-global aggregation layer and local guided aggregation layer to capture local and the whole-image cost dependencies respectively.
-Based on GANet [70], DSMNet [71] constructs a robust “domain normalization” method to overcome the challenges in cross-domain generalization.
 CoEx [2] proposes a novel approach called Guided Cost volume Excitation (GCE), which leverages image guidance to construct a simple channel excitation of the cost volume.
 IGEV-Stereo [55] leverages an iterative geometry encoding volume to capture local and non-local geometry information, outperforming existing methods on KITTI benchmarks and achieving cross-dataset generalization and high inference efficiency.
-Besides, as a potential solution to reduce the need for humans to design a good volumetric method, Neural Architecture Search (NAS) [20, 34, 4] has been employed in stereo matching [10].
 However, these CVM-Conv3D methods still suffer from low time efficiency and high memory requirements, which are far from real-time inference even on server GPUs.
@@
 Furthermore, FADNet++ [48] proposes an innovative approach to efficient disparity refinement using residual learning [17] in a coarse-to-fine manner.
-AutoDispNet [41] applied neural architecture search to automatically design stereo matching network structures.
+AutoDispNet [36] applied neural architecture search to automatically design stereo-matching network structures.
 More recently, Croco-Stereo [51] shows that large-scale pre-training can be successful for stereo matching through well-adapted pretext tasks.
@@
 3 OpenStereo
-In recent years, there has been a proliferation of new frameworks and evaluation datasets for stereo matching.
+In recent years, there has been a proliferation of new frameworks and evaluation datasets for stereo-matching.
 However, the lack of a unified and fair evaluation platform in this field is a significant issue that cannot be ignored.
 To address this challenge and promote academic research and practical application we have developed OpenStereo, a pyTorch-based [33] toolbox that provides a reliable and standardized evaluation framework for stereo matching.
-Figure 2: The design principles of proposed codebase OpenStereo.
 3.1 Design Principles of OpenStereo
@@
 However, OpenStereo provides a solution to this problem by supporting all of the aforementioned frameworks.
-With OpenStereo, researchers and practitioners can easily compare and evaluate different stereo matching models under a standardized evaluation protocol.
+With OpenStereo, researchers and practitioners can easily compare and evaluate different stereo-matching models under a standardized evaluation protocol.
 Support for various evaluation datasets.
-OpenStereo is a comprehensive tool that not only supports synthetic stereo datasets such as SceneFlow [35], but also five real-world datasets: KITTI2012 [14], KITTI2015 [36], ETH3D [45], Middlebury [43], and DrivingStereo [66].
+OpenStereo is a comprehensive tool that not only supports synthetic stereo datasets such as SceneFlow [30], but also four real-world datasets: KITTI2012 [12], KITTI2015 [31], ETH3D [40], and Middlebury [39].
 We introduce a suite of bespoke functions, meticulously crafted for each dataset, encompassing everything from initial data preprocessing to the final stages of evaluation.
@@
 As shown in Figure 3, the performance metrics we achieved, in most cases, surpass those reported in their original publications.
-3.2 Main Modules
-Technically, we follow the design of most PyTorch deep learning projects and divide OpenStereo into three modules, data, modeling, and evaluation.
-Data module includes three key components: data loader, which is responsible for loading and preprocessing the data; data sampler, which is used to select a subset of data for each iteration; and data transform, which applies various transformations to the data to increase the diversity and complexity of the training set.
-Modeling module is built upon the BaseModel base class, which predefines many behaviors of the stereo matching during both training and testing phases.
-This module consists of four essential components, namely the Feature Extraction, Cost Calculation, Cost Aggregation, and Disparity Processor, which are critical to current stereo matching algorithms.
-Evaluation module is utilized to assess the performance of the obtained model.
-Recognizing that different datasets require distinct evaluation protocols, we have incorporated these varied protocols directly into OpenStereo, significantly simplifying the evaluation process.
 4 Revisit Deep Stereo Matching
-4.1 Recheck on Previous Methods
-Figure 3: Quantitative evaluation on the SceneFlow [35] test and KITTI2015 [36] leadboard.
+Figure 3: Quantitative evaluation on the SceneFlow [30] and KITTI2015 [31] leadboard.
 For each model, the specific category on the SceneFlow used is consistent with the original paper.
 Underline refers to evaluation in the non-occluded regions only STTR [28].
+4.1 Datasets and Evaluation Metrics
+SceneFlow [30] is a synthetic stereo collection, counting 35,454/4,370 image pairs for training/testing, respectively.
+The data is split into two categories: cleanpass and finalpass.
+Cleanpass refers to the synthetic images that are generated by clean renderings without post-processing, while finalpass images are produced with photorealistic settings such as motion blur, defocus blur, and noise.
+The end point error (EPE) is utilized as the evaluation metric.
+KITTI.
+KITTI 2012 [12] and KITTI 2015 [31] are datasets that captured from real-world scenes.
+KITTI 2012 contains 194/195 image pairs for training/testing, while KITTI 2015 provides 200/200 image pairs for training and testing.
+All KITTI datasets provide sparse ground-truth disparities guided by the LiDAR system.
+For evaluations, we calculate EPE and the percentage of pixels with EPE larger than 3 pixels in all (D1-all) regions.
+All two KITTI datasets are also used for cross-domain generalization performance evaluation.
+Middlebury 2014 [39] has two sets of 15 image pairs for both training and testing, respectively.
+The stereo images are captured from indoor scenes, with 3 resolution options, which are full, half, and quarter.
+In evaluations, only the training image pairs with half resolution will be used to evaluate the cross-domain generalization performance, and EPE and 2px Error metric are reported.
+ETH3D [40] has stereo images collected from both indoor and outdoor environments with grayscale formate.
+It comes with 27 and 20 image pairs for training and testing, respectively, and a sparsely labeled ground truth for reference.
+In evaluation, only the training set is used to report the cross-domain generalization performance.
+And we adopt the EPE and 1px Error as the metric.
+4.2 Evaluation of Prior Work
 For benchmarking, it is critical to ensure that the results are reliable and trustworthy.
 To achieve this, we conducted our experiments on SceneFlow [30] and KITTI2015 [31] datasets.
+Regarding the KITTI2015 dataset, submission constraints led us to limit our leaderboard contributions to reproductions of the widely recognized PSMNet [5] and the latest SOTA IGEV [55].
+(More details in the Supplementary Material).
 As shown in Figure 3, the reproduced performances of OpenStereo are better than the results reported by the original papers.
-Regarding the KITTI2015 dataset, submission constraints led us to limit our leaderboard contributions to reproductions of the widely recognized PSMNet [6] and the latest state-of-the-art IGEV [60].
-OpenStereo is designed to offer the research community in stereo matching a standardized, comprehensive platform for method assessment.
-This facility enables meaningful and comparative analyses across various models, thereby serving as an valuable tool for the innovation and scrutiny of emerging algorithms.
-4.2 Denoising Stereo Matching
-With the support of OpenStereo, a comprehensive reevaluation of various stereo-matching methods is conducted, including data augmentation, feature extraction, cost construction, disparity prediction, and refinement.
-Our ablation studies have revealed some new insights.
-Figure 4: The visualization of stereo images and disparity with different data augmentation.
-The blue box represents the area where random cropping occurs during training.
-Notably, when the views are horizontally flipped, the disparity are multiplied by 11-1- 1, making the disparity map appear pure white.
-Table 1: The model undergoes training and evaluation on the finalpass of the SceneFlow [35] training and test sets.
-KITTI2015 [36] training set, consisting of 200 images, is only employed to evaluate the generalizability of models.
-RC stands for RandomCrop.
-HFlip denotes both images of a stereo and disparity are horizontally flipped.
-HSFlip horizontally flips both images in the stereo pair and then swaps them.
-VFlip involves vertically flipping both images in the stereo pair along with the disparity, inverting their top-bottom orientation.
-CES represents ColorAug, Erase, and Scale.
-Settings used in our final model are underlined.
+This is primarily due to the normalization applied to the input stereo images by the OpenStereo framework.
+(All other settings remain consistent with the original paper.)
+4.3 Necessity of Comprehensive Ablation Study
+In the evolving landscape of stereo-matching, comprehensive ablation studies play a pivotal role in deciphering the effectiveness of different components and strategies.
+A thorough ablation study goes beyond mere performance metrics; it uncovers the underlying mechanics of different algorithms, revealing their strengths and weaknesses in various scenarios.
+For instance, different data augmentation techniques may yield contrasting effects on the model’s ability to match stereo images accurately.
+Similarly, the impact of various backbones, cost volume configurations, and disparity regression methods on the overall performance can be profound.
+Understanding the specific contributions of each component is crucial for building more efficient and effective stereo-matching systems.
 Data Augmentation LR_scheduler SceneFlow KITTI15
@@
 […]
-Table 2: Ablation study of different backbones on the finalpass of SceneFlow test datasets [35].
-Flops and Params represent the computational complexity and parameters within the whole model, respectively.
+Table 1: Ablation study on Data Augmentation and LR_scheduler Selection.
+KITTI2015 [31] training set, consisting of 200 images, is only employed to evaluate the generalizability of models.
+RC stands for RandomCrop [23].
+HFlip [23] denotes both images of a stereo and disparity are horizontally flipped.
+HSFlip [23] horizontally flips both images in the stereo pair and then swaps them.
+VFlip [23] involves vertically flipping both images in the stereo pair along with the disparity, inverting their top-bottom orientation.
+CES represents ColorAug [23], Erase [69], and Scale [43].
+Settings used in our final model are underlined.
+4.4 Denoising Stereo Matching
+With the support of OpenStereo, a comprehensive reevaluation of various stereo-matching methods is conducted, including data augmentation, feature extraction, cost construction, disparity prediction, and refinement.
+Our ablation studies have revealed some new insights.
+ImpletionDetails For the experiments of this subsection, we adopt the COEX [2] as the baseline model due to its speed.
+All experiments in this subsection are conducted on SceneFlow [30] datasets.
 FLOPs are calculated at a resolution of $544\times 960544 × 960$.
-Backbone Type Pretrain Flops Params EPE
-MobilenetV2 100 [42] CNN 70.58G 2.78M 0.7737
-MobilenetV2 100 [42] CNN 70.58G 2.78M 0.6155
-MobilenetV2 100 [42] CNN 70.58G 2.78M 0.6155
-MobilenetV2 120d [42] CNN 85.93G 5.21M 0.5573
-EfficientNetV2 [48] CNN 157.52G 24.92M 0.5207
-RepViT [50] Trans. 101.35G 5.64M 0.5858
-MPViT [27] CNN&Trans. 283.35G 13.33M 0.5113
+The total epochs is set to 50.
+Figure 4: The visualization of stereo images and disparity with different data augmentation.
+The blue box represents the area where random cropping occurs during training.
+Notably, when the views are horizontally flipped, the disparity is multiplied by 11-1- 1, making the disparity map appear pure white.
 4.4.1 LR_scheduler and Data Augmentation
@@
 Visualizations of these augmentations are provided in Figure 4 for further insight.
-As shown in Table 1, most data augmentation strategies, except for the combined use of HSFlip and random crop, lead to a decline in the model’s EPE metric of SceneFlow.
+Most data augmentation strategies, except for the combined use of HSFlip and random crop, lead to a decline in the model’s EPE metric of SceneFlow.
 This is because stereo matching involves pixel-level matching, and these data augmentations (color augmentation and spatial transform) can affect the alignment of pixels.
@@
 These findings underscore the importance of selecting appropriate data augmentation methods to enhance model accuracy and robustness.
+Backbone Type Pretrain Flops Params EPE
+MobilenetV2 [37] CNN 70.58G 2.78M 0.77
+MobilenetV2 [37] CNN 70.58G 2.78M 0.62
+MobilenetV2 [37] CNN 70.58G 2.78M 0.62
+MobilenetV2 [37] CNN 85.93G 5.21M 0.56
+EfficientNetV2 [44] CNN 157.52G 24.92M 0.52
+RepViT [46] Trans. 101.35G 5.64M 0.59
+MPViT [24] CNN&Trans. 283.35G 13.33M 0.51
+Table 2: Ablation study on Backbones Selection. MobilenetV2 refers to MobilenetV2 120d. Bold: Best.
 4.4.2 Feature Extraction
@@
 To the best of our knowledge, our work is the first to explore the transformer-based feature extraction and the combination of CNN and transformer feature extraction for stereo matching.
-Table 3: Ablation study of cost volume on the finalpass of SceneFlow test datasets [35].
-Gwc represents Group-wise correlation volume [18].
-Cat stands for Concatenation volume [6].
-G8-C16, G16-C24, and G32-C48 combine Gwc volume and Cat volume [18].
-Channel and Dims represent the channel and dimensions of the cost volume, respectively.
-FLOPs are calculated at a resolution of $544\times 960544 × 960$.
 Cost Volume Dims Channel Flops Params EPE
@@
 G32-C48 4D 80 3239.17G 39.37M 0.60
+Table 3: Ablation study on Cost Construction.
+Gwc represents Group-wise correlation volume [15].
+Cat stands for Concatenation volume [5].
+G8-C16, G16-C24, and G32-C48 combine Gwc volume and Cat volume [15].
+Channel and Dims represent the channel and dimensions of the cost volume, respectively.
+Bold: Best.
 4.4.3 Cost Construction
@@
 While deeper and combined volumes reduce the EPE, the gains might be marginal compared to the significant increase in computational requirements, raising questions about the practicality of these approaches in resource-constrained environments.
-Table 4: Ablation study of disparity regression and refinement on the finalpass of SceneFlow test datasets [35]. ArgMin refers to Differentiable ArMin. Context stands for ContextUpasmple [33, 60].
 Regression Refinement Flops Params EPE
@@
 ArgMin ConvGRU [29, 55] 3023.88G 12.51M 0.46
+Table 4: Ablation study on Disparity Regression and Refinement. ArgMin refers to Differentiable ArMin. Context stands for ContextUpasmple [29, 55]. Bold: Best.
+Figure 5: Overview of our proposed StereoBase. GwcVolume represents Group-wise correlation volume [15]. CGEV refers to Combined Geometry Encoding Volume [55].
 4.4.4 Disparity Regression and Refinement
@@
 Combining these methods further reduces EPE to 0.69, indicating marginal benefits from their integration.
-However, employing ConvGRU refinement substantially improves EPE 0.46, albeit at a significant cost in computational complexity (3023.88 Gflops) and model size (12.51M).
+However, ConvGRU refinement substantially improves EPE to 0.46, albeit at a significant cost in computational complexity (3023.88 Gflops) and model size (12.51M).
 This highlights a trade-off between accuracy improvements and increased computational demands.
-4.3 Analysis and Discussion
-4.3.1 Necessity of Comprehensive Ablation Study
-In the evolving landscape of deep stereo matching, comprehensive ablation studies play a pivotal role in deciphering the effectiveness of different components and strategies.
-A thorough ablation study goes beyond mere performance metrics; it uncovers the underlying mechanics of different algorithms, revealing their strengths and weaknesses in various scenarios.
-For instance, different data augmentation techniques may yield contrasting effects on the model’s ability to match stereo images accurately.
-Similarly, the impact of varying backbones, cost volume configurations, and disparity regression methods on the overall performance can be profound.
-Understanding the specific contributions of each component is crucial for building more efficient and effective stereo-matching systems.
-4.3.2 Necessity of A Strong Baseline
+5 A Strong Pipeline: StereoBase
 A strong baseline in deep stereo-matching research is critical for several key reasons.
@@
 This is crucial for maintaining consistency and validity in comparative analyses.
-In summary, a strong baseline is essential for meaningful advancements in deep stereo matching, ensuring that new developments are substantial, accurately assessed, and broadly applicable.
-5 A Strong Pipeline: StereoBase
+Method SceneFlow KITTI 2012 KITTI 2015
+EPE 3-noc 3-all D1-bg D1-fg D1-all
+PSMNet [5] 1.09 1.49 1.89 1.86 4.62 2.32
+GwcNet [15] 0.76 1.32 1.70 1.74 3.93 2.11
+CFNet [42] 1.04 1.23 1.58 1.54 3.56 1.88
+CascadeStereo [13] 0.72 - - 1.59 4.03 2.00
+IGEV-Stereo [55] 0.47 1.12 1.44 1.38 2.67 1.59
+GANet+ADL [57] 0.50 0.98 1.29 1.38 2.38 1.55
+NMRF-Stereo [14] 0.45 1.01 1.35 1.28 3.13 1.59
+Selective-IGEV [49] 0.44 1.07 1.38 1.33 2.61 1.55
+MoCha-Stereo [7] 0.41 1.06 1.35 1.36 2.43 1.53
+StereoBase (Ours) 0.34 1.00 1.26 1.28 2.26 1.44
+Table 5: Results on SceneFlow [30], KITTI 2012 [12], and KITTI 2015 [31] leaderboard. Bold: Best.
+Figure 6: Visualization results on KITTI2015 dataset.
 5.1 Pipeline
@@
 Hourglass networks [55] were implemented for cost aggregation, while convGRU [55] strategies were applied for the final disparity regression.
-Figure 5: Overview of our proposed StereoBase. GwcVolume represents Group-wise correlation volume [18]. CGEV refers to Combined Geometry Encoding Volume [60].
-Table 5: Results on SceneFlow [35], KITTI 2012 [14], KITTI 2015 [36] leaderboard, and DrivingStereo [66].
-All results on DrivingStereo [66] are derived using the OpenStereo.
-Underline refers to evaluation in the non-occluded regions only STTR [31].
+Method K12 K15 Midd ETH3D
+STTR [28] 49.72 40.26 OOM 38.89
+PSMNet [5] 30.51 32.15 33.53 18.02
+CFNet [42] 13.64 12.09 23.91 7.67
+AANet [56] 7.23 7.72 22.45 18.77
+M2D [41] 18.34 21.21 34.04 13.89
+M3D [41] 25.14 22.44 29.32 13.71
+GwcNet [15] 23.05 25.19 29.87 14.54
+COEX [2] 12.08 11.01 25.17 11.43
+FADNet++ [48] 11.31 13.23 24.07 22.48
+CStereo [13] 11.86 12.06 27.39 11.62
+IGEV [55] 4.88 5.16 8.47 3.53
+StereoBase (Ours) 4.85 5.35 9.76 3.12
+Table 6: Cross-domain evaluation on Middlebury, ETH3D, and KITTI all training sets.
+All methods are only trained on the Scene Flow dataset.
+Middlebury is tested on half-resolution.
+The model with indicates the implementation of OpenStereo.
 Bold: Best.
-Method SceneFlow KITTI 2012 KITTI 2015 DrivingStereo
-EPE 3-noc 3-all D1-bg D1-fg D1-all EPE D1-all
-STTR [31] 0.43 - - 1.70 3.61 2.01 OOM OOM
-PSMNet [6] 1.09 1.49 1.89 1.86 4.62 2.32 1.19 2.26
-GwcNet [18] 0.76 1.32 1.70 1.74 3.93 2.11 0.99 1.36
-CFNet [47] 1.04 1.23 1.58 1.54 3.56 1.88 0.98 1.46
-AANet [61] 0.87 1.91 2.42 1.65 3.96 2.03 2.91 15.16
-Mobilestereo-3D [46] 0.80 - - 1.75 3.87 2.10 1.06 1.61
-COEX  [2] 0.68 1.55 1.93 1.74 3.41 2.02 1.34 2.70
-FADNet++  [52] 0.76 - - 1.99 3.18 2.19 1.44 5.15
-CascadeStereo [15] 0.72 - - 1.59 4.03 2.00 1.31 2.84
-IGEV-Stereo [60] 0.47 1.12 1.44 1.38 2.67 1.59 1.06 1.50
-GANet+ADL [63] 0.50 0.98 1.29 1.38 2.38 1.55 - -
-NMRF-Stereo [17] 0.45 1.01 1.35 1.28 3.13 1.59 - -
-Selective-IGEV [53] 0.44 1.07 1.38 1.33 2.61 1.55 - -
-MoCha-Stereo [8] 0.41 1.06 1.35 1.36 2.43 1.53 - -
-StereoBase(Ours) 0.34 1.00 1.26 1.28 2.26 1.44 1.15 2.19
 5.2 Comparison with State-of-the-art Methods
-In our comprehensive evaluation, we benchmarked StereoBase against current state-of-the-art methods on SceneFlow [35], KITTI2012 [14], 2015 [36], and DrivingStereo [66].
+Implementation details Training is driven by the smooth L1 [5] loss function.
+We utilize the AdamW optimizer in conjunction with the OneCycle learning rate scheduler.
+For SceneFlow [30], the total training epochs is set to 90.
+On the KITTI dataset, we finetune the pre-trained model using a combined dataset of KITTI2012 [12] and 2015 [31] training image pairs.
+The total epoch is 500.
+(More implementation details in the Supplementary Material).
+In our comprehensive evaluation, we benchmarked StereoBase against current state-of-the-art methods on SceneFlow [30], KITTI2012 [12], and KITTI2015 [31].
 On the SceneFlow [30] test set, we achieve a new SOTA EPE of 0.34.
@@
 This evaluation further validates the adaptability and potential of StereoBase in handling diverse and challenging stereo vision tasks.
-Table 6: Cross-domain evaluation on Middlebury, ETH3D, and KITTI all training sets.
-All methods are only trained on the Scene Flow datatest.
-Middlebury is tested on half-resolution.
-The model with indicates the implementation of OpenStereo.
-Bold: Best.
-Method KITTI2012 KITTI2015 Middlebury ETH3D
-D1-all(%) EPE D1-all(%) EPE bad 2.0(%) bad 1.0(%)
-STTR 49.72 6.80 40.26 6.16 OOM 38.89
-PSMNet 30.51 4.68 32.15 5.99 33.53 18.02
-CFNet 13.64 2.27 12.09 2.89 23.91 7.67
-AANet 7.23 1.27 7.72 1.41 22.45 18.77
-Mobilestereo-2D 18.34 2.45 21.21 2.78 34.04 13.89
-Mobilestereo-3D 18.96 2.79 19.69 3.40 29.32 13.71
-GwcNet 23.05 2.76 25.19 3.58 29.87 14.54
-COEX 12.08 1.80 11.00 2.48 25.17 11.43
-FADNet++ 11.31 1.77 13.23 2.97 24.17 25.53
-CascadeStereo 11.83 1.83 12.03 2.69 27.27 11.68
-IGEV 4.88 0.98 5.16 1.18 8.47 3.53
-StereoBase 4.85 0.99 5.35 1.18 9.76 3.12
-Figure 6: Visualization results on KITTI2015 dataset.
 6 Conclusion
@@
 Through OpenStereo and StereoBase, we aim to contribute a substantial and versatile resource to the stereo-matching community, fostering innovation and facilitating more effective and efficient research.
-Acknowledgements.
-This work was supported by the National Natural Science Foundation of China under Grant 62373356 and the Open Projects Program of State Key Laboratory of Multimodal Artificial Intelligence Systems.